### v2

# Valid F-screening in linear regression

Abstract
[…]
1 Introduction
Not all data analyses yield interesting results, and many uninteresting findings are left in the “file drawer” (Rosenthal, 1979) rather than published. Furthermore, when papers are published, analyses yielding null results are often omitted. Such selective reporting has been empirically observed in the biomedical literature (Chan et al., 2004; Ioannidis, 2005).
Why is selective reporting a problem? Standard statistical guarantees, such as Type 1 error control for hypothesis tests and nominal coverage of confidence intervals, only hold when results are reported regardless of whether they are “interesting.”
In this work, we consider a particular and pervasive form of selective reporting that arises in the context of a regression model of the form
$Y=\beta_{0}+X_{1}\beta_{1}+\ldots+X_{p}\beta_{p}+\epsilon,\;\;\;\epsilon\sim\mathcal{N}(0,\sigma^{2}).$ (1)
Given $n$ independent observations from (1), suppose that the data analyst first tests
$H_{0}^{1:p}:\beta_{1}=\beta_{2}=\ldots=\beta_{p}=0$ (2)
at level $\alpha_{0}\in(0,1)$ in order to determine whether the model in (1) merits further study. If they fail to reject $H_{0}^{1:p}$, then (1) receives no further consideration, and the fact that (1) was considered may not even be reported. On the other hand, if (and only if) they reject $H_{0}^{1:p}$, they then proceed to test $H_{0}^{M}:\beta_{M}=0$ for fixed $M\subseteq\left\{1,\ldots,p\right\}$, where $\beta_{M}:=\{\beta_{j}:j\in M\}.$ Because $H_{0}^{1:p}$ is typically tested by means of an $F$-statistic, we refer to this procedure—summarized in Box 1—as F-screening. If conducted, the test in Step 2 of Box 1 is sometimes referred to as a post hoc test.
Box 1: F-screening in the model (1). Step 1: Test $H_{0}^{1:p}$ as given in (2) at level $\alpha_{0}$, using an $F$-test. Step 2: Only if $H_{0}^{1:p}$ is rejected, conduct inference on $\beta_{M}$ for fixed $M\subseteq\{1,\ldots,p\}$.
To see that conducting “standard” inference in Step 2 of F-screening will not provide the usual statistical guarantees, consider the extreme setting where $p=1$ and $\beta_{1}=0$. In this case, $H_{0}^{1:p}\equiv H_{0}^{1}$. Suppose we carry out F-screening in Box 1 using the “standard” $t$-test for $H_{0}^{1}$ in Step 2. Then for fixed $\alpha_{0}$, the tests in Steps 1 and 2 are identical, and so we will report a result in Step 2 only if $H_{0}^{1}$ is rejected. Thus the Type 1 error rate among reported results equals one! The left plot in Fig. 5 illustrates the inflation of Type 1 error that results from conducting “standard” inference in Step 2 in a setting with $p=10$ and $n=100$.
As pointed out by Olshen (1973), conducting a test of $H_{0}^{M}$ that accounts for rejection of $H_{0}^{1:p}$ is distinct from the task of correcting multiple tests of the form $H_{0}^{M}:\beta_{M}=0$ for multiple comparisons. The latter has been very well-explored in the literature: see Dunn (1961); Šidák (1967); Tukey (1949); Scheffé (1953). For instance, if our interest in Step 2 of Box 1 lies in $H_{0}^{j}:\beta_{j}=0$ for $j\in\{1,\ldots,p\}$, then we could consider testing each null hypothesis using a standard test, and applying a Bonferroni correction to account for the fact that $p$ tests were conducted. However, this will not address the fact that F-screening was performed: namely, that we test $\{H_{0}^{j}\}_{j=1}^{p}$ if and only if we have rejected $H_{0}^{1:p}$. In Appendix A, we elaborate on how the guarantees provided by multiple testing corrections—such as family-wise error rate control—differ from those achieved by our proposed method. Further, we describe a closed-testing procedure that parallels F-screening and show that its guarantees are quite different. Fig. A.1 demonstrates empirically that existing approaches fail to provide valid inference conditional on the rejection of $H_{0}^{1:p}$.
Our goal of conducting inference on $\beta_{M}$ that is valid conditional on the fact that we rejected $H_{0}^{1:p}$ falls under the umbrella of selective inference or post-selection inference; we will view rejection of $H_{0}^{1:p}$ as the “selection event.” The selective inference framework was laid out in Fithian et al. (2017) and further explored in a variety of settings (Taylor & Tibshirani, 2015; Lee et al., 2016; Zhang et al., 2019; Chen & Bien, 2020; Jewell et al., 2022; Chen & Gao, 2024; Chen et al., 2023; Chen & Witten, 2023; Gao et al., 2024; Neufeld et al., 2022, among others). Our goal also relates to the discussion of selective reporting of confidence intervals in Benjamini & Yekutieli (2005). Most related to our work is Heller et al. (2019), who considered inference following aggregate testing for a coefficient that follows a multivariate normal distribution with known covariance. We instead consider inference on the regression coefficients in Step 2 of Box 1.
Our contributions are as follows:
  1. 1.
    We provide a valid end-to-end approach for conducting F-screening in the sense of Box 1. That is, we enable reporting of all of the standard multiple linear regression outputs—such as p-values, point estimates, and confidence intervals—conditional on rejection of $H_{0}^{1:p}$ in Step 1; we will refer to them as selective p-values, point estimates, and confidence intervals.
  2. 2.
    We establish consistency and selective Type 1 error control (the Type 1 error conditional on rejection of $H_{0}^{1:p}$) of the test resulting from our selective p-value.
  3. 3.
    We characterize the Fisher information about $\beta_{M}$ resulting from our procedure, and show that it compares favorably to that of a procedure that relies on sample splitting.
  4. 4.
    We show that the selective p-values can be computed using only the standard multiple linear regression outputs without access to the raw data. This is particularly critical because, while it is firmly established that selective reporting is common practice in the published literature (Chan et al., 2004), the raw data is typically not available to the statistician who wishes to correct for it.
  5. 5.
    We specialize our approach to the setting of analysis of variance (ANOVA), in which the F-screening procedure in Box 1, with Step 2 conducted using a “standard” approach that does not account for Step 1, is particularly pervasive (see, e.g., Fig. 1).
Figure 1: A subset of Table 1 from Keil et al. (2023). The $n=826$ participants in the dataset are split into three age groups: under 65 years old, between 65 and 85 years old, and over 85 years old. Each row in the table corresponds to a continuous variable of interest, with means and standard deviations reported for each age group. For each of these continuous variables, ANOVA is conducted to test for a difference in means between the three age groups (corresponding to Step 1 of Box 1). The associated p-value is reported in the “P Value” column. The last column holds p-values for “post hoc tests” which are typically conducted only if the p-value associated with the ANOVA test is small (corresponding to Step 2 of Box 1). In this case, the p-values in the last column would be invalid since they do not account for the fact that they were only computed because $H_{0}^{1:p}$ (2) was rejected.
Addressing the inferential problems posed by F-screening is especially important from an applied perspective. A theoretical statistician unaccustomed to applied collaboration might believe that F-screening is an unusual procedure unlikely to be conducted in practice, or that conducting Step 2 of Box 1 without accounting for Step 1 is “obviously wrong” and no data analyst would consider it. The theoretical statistician would be mistaken on both counts. Statistical references that encourage F-screening with “standard” inference in Step 2 as part of a routine data analysis abound, and are authored by many of the luminaries in the field. [^1] Fisher (1935, §24) writes: “If the yields of the different varieties in the experiment fail to satisfy the test of significance they will not often need to be considered further . . . If, however . . . the null hypothesis has been falsified . . . we shall thereafter proceed to interpret the differences between the varietal yields . . . and shall be concerned to know with what precision these different yields have been evaluated.” A few years later, Lindquist (1940, page 98) writes “The student may well ask why it was necessary to apply the $F$-test at all, if we were subsequently going to test the individual differences anyway. The answer is that the $F$-test tells us whether it is worthwhile to test the individual difference at all. If the $F$-test had not proven significant, we would have known at once that all observed differences in . . . means could be due to chance alone. In that case it would not only have been unnecessary, but decidedly improper to apply the $t$-test to individual differences.” The abstract of Scheffé (1953) poses the following question: “Under the usual assumptions, if the conventional $F$-test of $H:\mu_{1}=\mu$ . . . rejects $H$, what further inference are valid about the contrasts $H_{i}$?” Scheffé (1953) later remarks that “The problem of making further inferences about the contrasts, arising when the $F$-test rejects $H$, has been considered by various writers, including Fisher (1935), Newman (1939), Tukey (1949), Nandi (1951).” As pointed out in Olshen (1973), to the extent that these references consider selection, they focus on the multiplicity issues that arise from testing $H_{0}^{j}:\beta_{j}=0$, $j=1,\ldots,p$, rather than on the challenge that arises from inference conditional on rejecting $H_{0}^{1:p}$. See Appendix A for simulation results demonstrating that multiple testing corrections do not correct for selection in the F-screening procedure. Thus, it is hardly surprising that in the scientific literature, conducting Step 2 of Box 1 with a “standard” approach that does not account for having conducted Step 1 is often viewed as the correct analysis. [^2] As one of many possible examples, Webb et al. (2024) describe their analysis as follows: “A one-way ANOVA revealed that amygdala reactivity was not significantly different by trajectory groups . . . Therefore, additional analyses testing whether greenspace was associated with trajectory assignment via amygdala reactivity were not conducted.” In clinical journals, F-screening is a routine part of “Table 1,” in which patient characteristics are summarized; an example is shown in Fig. 1.
Finally, we note that F-screening is likely conducted much more frequently than the published literature suggests, both because (i) it is typically unknown whether an analysis would have been included in a published paper if $H_{0}^{1:p}$ had not been rejected, and (ii) an analysis for which $H_{0}^{1:p}$ is not rejected may not be published at all. Given that F-screening is both commonplace and problematic if inference in Step 2 does not account for Step 1, it came as a surprise to us that—to the best of our knowledge—no statistical solution has yet been proposed for valid inference in Step 2. The rest of this paper fills this gap.
The paper is organized as follows. In Section 2, we propose a conditional selective inference approach for valid inference in Step 2 of Box 1. We establish its theoretical properties and derive an expression for the leftover Fisher information after rejection of $H_{0}^{1:p}$ (2) in Step 1. We also propose an estimator of $\sigma^{2}$ that is unbiased and consistent conditional on rejection of $H_{0}^{1:p}$, and that yields valid p-values. Section 3 shows that it is possible to compute selective p-values using the results of a “standard” (non-selective) analysis in Step 2, without having access to the raw data. Section 4 specializes the results in Sections 2 and 3 to the case of ANOVA. Simulation results are in Section 5 (and Appendix C), and re-analyses of published findings in Section 6. The discussion is in Section 7. The main text focuses on developing valid selective p-values; confidence intervals and point estimates are deferred to Appendix B. Appendix D contains mathematical details for Figures 2 and 3. Proofs of all results are given in Appendix E.
2 Developing valid p-values in Step 2 of F-screening
2.1 Preliminaries
Let $P_{A}:=A(A^{\top}A)^{-1}A^{\top}$ denote the projection onto the column space of a matrix $A$ with full column rank, and let $A_{-S}$ denote the matrix $A$ with the columns indexed by $S\subseteq\{1,\ldots,p\}$ removed. Furthermore, let $\mathbbm{1}_{n}$ denote the $n$-dimensional column vector of 1’s, and $I_{n}$ the $n\times n$ identity matrix. This section restates some standard results in the theory of linear models (a more thorough treatment can be found elsewhere, e.g., Seber & Lee, 2003).
Consider $n$ independent observations from the model in (1). We assume that the intercept is not of scientific interest, and so both $H_{0}^{1:p}$ and $H_{0}^{M}$ will operate on centered data. To do this, we let $U\in\mathbb{R}^{n\times(n-1)}$ be any matrix satisfying both $U^{\top}U=I_{n-1}$ and $UU^{\top}=I_{n}-P_{\mathbbm{1}_{n}}$, and then with slight abuse of notation we redefine $Y:=U^{\top}Y$, $X_{1}:=U^{\top}X_{1}$, and so on. Thus, in what follows, we will treat $\mathbb{R}^{n-1}$ as the ambient space and dispense with the intercept. For now, we assume that $\sigma^{2}$ is known (but that $F$-statistics are used nonetheless); we address the setting of unknown variance in Section 2.4.
First, consider the null hypothesis $H_{0}^{1:p}$ given in (2). Under $H_{0}^{1:p}$, the $F$-statistic defined as
${F_{H_{0}^{1:p}}}:=\frac{n-p-1}{p}\cdot\frac{Y^{\top}P_{X}Y}{Y^{\top}(I_{n-1}-P_{X})Y}$ (3)
follows an $F_{p,n-p-1}$ distribution (that is, an $F$ distribution with $p$ and np1n-p-1 degrees of freedom). Next, fix $H_{0}^{M}$ by letting $M\subseteq\{1,\ldots,p\}$ denote a set of indices with cardinality $\lvert M\rvert=m$, and let $\beta_{M}$ denote the subset of coefficients indexed by $M$ (in the simplest case, $M=\{j\}$ for fixed $j\in\left\{1,\ldots,p\right\}$). Under the null hypothesis $H_{0}^{M}$, the $F$-statistic
$F_{H_{0}^{M}}:=\frac{n-p-1}{m}\cdot\frac{Y^{\top}(P_{X}-P_{X_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}$ (4)
follows an $F_{m,n-p-1}$ distribution. Thus, for a realization $y$ of the random variable $Y$, we can compute the p-value for $H_{0}^{M}:\beta_{M}=0$ as
$p_{H_{0}^{M}}(y)$ $:=\Pr_{\beta_{M}=0}\left(\frac{Y^{\top}(P_{X}-P_{X_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\right)$
$=1-F_{m,n-p-1}\left(\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\cdot\frac{(n-p-1)}{m}\right),$ (5)
where $F_{m,n-p-1}\left(\cdot\right)$ is the cumulative distribution function of a random variable with an $F_{m,n-p-1}$ distribution. This is the “standard” test of $H_{0}^{M}:\beta_{M}=0$, i.e., it does not account for Step 1 of Box 1.
Fig. 2 displays the rejection regions corresponding to a test of $H_{0}^{1:p}$ using (3) and a test of $H_{0}^{M}$ using (4) in a simple setting with $n=3$, $p=2$, and $X$ orthogonal. The geometry of these two events interplay in non-trivial ways, and failing to account for the first when evaluating the second can yield invalid inference. By contrast, the selective test that we will develop in the next section will take the sequential nature of the analysis into account, and it will adjust the geometry of the second test accordingly.
Figure 2: We consider testing $H_{0}^{1:2}:\beta_{1}=\beta_{2}=0$ using $F_{H_{0}^{1:2}}$ (3) and $H_{0}^{1}:\beta_{1}=0$ using $F_{H_{0}^{1}}$ (4) in the model $Y=X\beta+\epsilon$, where $X=\begin{bmatrix}1&0\\ 0&1\\ 0&0\end{bmatrix}$. The red double cone corresponds to the rejection boundary of the test $H_{0}^{1:2}$, and the blue planes correspond to the rejection boundary of $H_{0}^{1}$. Thus data points that lie farther from the $Y_{3}$-axis than the double cone lead to rejection of $H_{0}^{1:2}$, and data points that lie farther from the $(Y_{2},Y_{3})$-plane than the two blue planes lead to rejection of $H_{0}^{1}$. See Appendix D for mathematical details.
2.2 A conditional selective inference approach
We test $H_{0}^{M}$ in Step 2 of Box 1 if and only if we reject $H_{0}^{1:p}$ (2) at level $\alpha_{0}\in(0,1)$: that is, if and only if $Y\in E_{1}$, where
$E_{1}:=\left\{Y:\frac{n-p-1}{p}\cdot\frac{Y^{\top}P_{X}Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq F^{-1}_{p,n-p-1}(1-\alpha_{0})\right\}.$ (6)
The left panel of Fig. 5 shows that conditional on $Y\in E_{1}$ (6), the p-values obtained using $F_{H_{0}^{M}}$ are far from uniform and indeed anti-conservative—i.e., a test based on $p_{H_{0}^{M}}$ (5) does not control the selective Type 1 error in Step 2 of Box 1. This is because $F_{H_{0}^{M}}$ (4) does not follow an $F_{m,n-p-1}$ distribution under $H_{0}^{M}$ when we condition on $Y\in E_{1}$ (6). Our goal is to develop an alternative to $p_{H_{0}^{M}}$ (5) that does lead to selective Type 1 error control.
To illustrate this idea, we consider in Fig. 3 a simple setting with $p=2$ and $\sigma^{2}$ known. The left-hand panel displays the rejection regions for $\chi^{2}$ tests of $H_{0}^{1:2}$ and $H_{0}^{1}$, respectively; this does not account for the sequential nature of the two tests in the context of F-screening. The right-hand panel displays the rejection region of a $\chi^{2}$ test of $H_{0}^{1:2}$, alongside the rejection region of a test of $H_{0}^{1}$ that conditions upon rejection of $H_{0}^{1:2}$.
Figure 3: We consider testing $H_{0}^{1:2}:\beta_{1}=\beta_{2}=0$ and $H_{0}^{1}:\beta_{1}=0$ in the model $Y=X\beta+\epsilon$ where $X=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ and we take the variance of $\epsilon$ to be known. In the left plot we use $\chi^{2}$-tests to test $H_{0}^{1:2}$ and $H_{0}^{1}$; this is in analogy to F-screening in Box 1 where Step 2 is conducted using a “standard” test. On the right, we use a $\chi^{2}$-test to test $H_{0}^{1:2}$, and then we use a test of $H_{0}^{1}$ in Step 2 that accounts for the rejection of $H_{0}^{1:2}$ in Step 1. This leads to a substantial change in the geometry of the rejection region of $H_{0}^{1}$. Mathematical details are in Appendix D.
To extend this idea to the setting of the paper, we make use of $F$-tests rather than $\chi^{2}$ tests. We modify $p_{H_{0}^{M}}$ in (5) to condition on the selection event in (6), leading to
$\Pr_{\beta_{M}=0}\left(\frac{Y^{\top}(P_{X}-P_{X_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\middle|\>Y\in E_{1}\right).$ (7)
Unfortunately, (7) is not a well-defined probability: it involves nuisance parameters $\left\{\beta_{j}\right\}_{j\notin M}$ not fixed under $H_{0}^{M}$, and thus it cannot be computed. To remedy this, we additionally condition on the event $E_{2}(y):=\left\{Y:Y^{\top}P_{{X}_{-M}}Y=y^{\top}P_{{X}_{-M}}y\right\};$ in effect, we hold fixed all the variance in $Y$ that is in the column space of $X_{-M}$. We then define
$E(y):=E_{1}\cap E_{2}(y).$ (8)
This leads to the p-value
${p_{H_{0}^{M}\mid E}}(y):=\Pr_{\beta_{M}=0}\left(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq\frac{y^{\top}(P_{X}-P_{{X}_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\middle|Y\in E(y)\right).$ (9)
The next result establishes that a test of $H_{0}^{M}$ based on ${p_{H_{0}^{M}\mid E}}(y)$ controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected.
Theorem 2.1.
A test of $H_{0}^{M}$ based on the p-value ${p_{H_{0}^{M}\mid E}}(y)$ (9) controls the selective Type 1 error, conditional on the event that $H_{0}^{1:p}$ was rejected. That is, for any $\alpha^{\prime}\in(0,1)$,
$\Pr_{\beta_{M}=0}\left({p_{H_{0}^{M}\mid E}}(Y)\leq\alpha^{\prime}\mid Y\in E_{1}\right)=\alpha^{\prime}.$
The left panel of Fig. 5 empirically corroborates this result. The next result shows that ${p_{H_{0}^{M}\mid E}}(y)$ can be easily approximated via Monte Carlo.
Proposition 2.1.
Define $Z\sim\chi_{n-p-1}^{2}$ and $W\sim\chi_{m}^{2}$, where $Z\perp W$ (i.e., $Z$ is independent of $W$). Then,
${p_{H_{0}^{M}\mid E}}(y)=\Pr\left(W/Z\geq r(y)\mid W-cZ\geq d(y)/\sigma^{2}\right),$ (10)
where
$c:=F^{-1}_{p,n-p-1}(1-\alpha_{0})\cdot\frac{p}{n-p-1},\;\;\;d(y):=-y^{\top}P_{X_{-M}}y,\;\;\;r(y):=\frac{y^{\top}(P_{X}-P_{{X}_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}.$ (11)
To summarize, a test based on ${p_{H_{0}^{M}\mid E}}(y)$ in (10) controls the selective Type 1 error for $H_{0}^{M}$ in the context of F-screening as described in Box 1 (Theorem 2.1). Furthermore, ${p_{H_{0}^{M}\mid E}}(y)$ can be easily approximated via Monte Carlo, provided that the error variance $\sigma^{2}$ is known (Proposition 2.1). Obtaining confidence intervals for $\beta_{1},\ldots,\beta_{p}$ that attain the nominal selective coverage (conditional on rejecting $H_{0}^{1:p}$), and correcting the corresponding point estimates for selection, are straightforward extensions of the ideas in this section. Details are provided in Appendix B.
We close this subsection with a result concerning the power of the test that results from using the p-value in (10).
Proposition 2.2.
Consider a fixed sequence of design matrices $X_{n}\in\mathbb{R}^{n\times p}$ and corresponding realizations of the linear model $Y_{n}=X_{n}\beta+\epsilon_{n}$ from (1). Suppose that the design matrices satisfy $\frac{1}{n}X_{n}^{\top}X_{n}\to Q$ for some positive definitive Q. Then the test of $H_{0}^{M}:\beta_{M}=0$ that uses the selective p-value in (10) is consistent under every alternative, conditional on $Y_{n}\in E_{1}$. That is, for any $\beta$ satisfying $\beta_{M}\neq 0$ and arbitrary $\alpha^{\prime}\in(0,1)$, $\lim_{n\to\infty}\;\Pr_{\beta}\left({p_{H_{0}^{M}\mid E}}\left(Y_{n}\right)\leq\alpha^{\prime}|Y_{n}\in E_{1}\right)=1.$
Remark 1.
In Proposition 2.2, the following quantities are fixed with respect to $n$: $p\geq 1$, $\beta\in\mathbb{R}^{p}$, $\sigma^{2}>0$, $M\subseteq\{1,\ldots,p\}$. In contrast, $X_{n}$ and $Y_{n}$ grow in length with $n$, and the p-value in (10) depends on $n$ through $Z\sim\chi^{2}_{n-p-1}$, as well as $c$, $d$, and $r$ in (11). In the remainder of the paper, for conciseness, we only include the subscript $n$ when we wish to emphasize dependence on $n$. We may also suppress the dependence of some quantities on $y$ (e.g., ${p_{H_{0}^{M}\mid E}}(y)$ in (10) may be written $p_{H_{0}^{M}\mid E}$}).
2.3 F-Screening and (Conditional) Fisher Information
We now consider the test of $H_{0}^{M}$ in Step 2 of Box 1 through the lens of Fisher information in a simplified setting in which $\sigma^{2}$ is known and $\beta_{0}=0$. Hence, we consider the model (1) without the intercept, and the F-tests in Steps 1 and 2 of Box 1 are replaced with chi-squared tests. The hypothesis of interest in Step 2 is $H_{0}^{1}:\beta_{1}=0$. We compare our proposed approach (specialized to this setting) to a sample splitting procedure.
In the sample splitting procedure, we partition the data into $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$ and $(X^{\mathrm{te}},Y^{\mathrm{te}})$, where a proportion $\rho\in(0,1)$ of the observations are assigned to the training set and the remaining observations are assigned to the test set such that $\rho n$ and $(1-\rho)n$ are integers. We test $H_{0}^{1:p}$ with $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$, and then $(X^{\mathrm{te}},Y^{\mathrm{te}})$ is used to test $H_{0}^{1}$ if and only if $H_{0}^{1:p}$ is rejected.
In the spirit of Fithian et al. (2017, Section 2.5), we derive expressions for: (i) the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on the event that $H_{0}^{1:p}$ is rejected with $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$; and (ii) the leftover Fisher information in $Y$ about $\beta_{1}$ conditional on the event that both the overall null $H_{0}^{1:p}$ is rejected using the full data $(X,Y)$ and $P_{X_{-1}}Y=P_{X_{-1}}y$ for some realization $y\in\mathbb{R}^{n}$.
Definition 1 (adapted from Fithian et al., 2017).
Let $\ell(\theta;Y\mid Y\in A)$ denote the conditional log-likelihood of $Y$ given $Y\in A$. The leftover Fisher information about the parameter $\theta$ in data $Y$ conditional on $Y\in A$ is $\mathcal{I}_{Y|Y\in A}(\theta\mid Y\in A):=-{\mathbb{E}}\left[\nabla^{2}\ell(\theta;Y\mid Y\in A)\mid Y\in A\right].$
Proposition 2.3.
  1. i.
    Define $c_{\chi}=\sigma^{2}(\chi^{2}_{p})^{-1}(1-\alpha_{0})$ and the set
    $R_{\rho}(X):=\{Y\in\mathbb{R}^{(\rho n)}:Y^{\top}P_{X}Y\geq c_{\chi}\}.$ (12)
    Treating $\beta_{2},\ldots,\beta_{p}$ as nuisance parameters, the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$ equals the (unconditional) Fisher information about $\beta_{1}$ in $Y^{\mathrm{te}}$. This quantity is $\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}\right)$.
  2. ii.
    For any $\tilde{y}\in\mathbb{R}^{n},$ define
    $R(\tilde{y};X):=\{Y\in\mathbb{R}^{n}:Y^{\top}P_{X}Y\geq c_{\chi},\;P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}\}$. (13)
    Treating $\beta_{2},\ldots,\beta_{p}$ as nuisance parameters, the leftover Fisher information in $Y$ about $\beta_{1}$ conditional on $Y\in R(\tilde{y};X)$ equals
    $\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))$
    $=\begin{cases}\tfrac{X_{1}^{\top}(I-P_{X_{-1}})X_{1}}{\sigma^{2}}&\text{if }\;\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}\\ \tfrac{X_{1}^{\top}(I-P_{X_{-1}})X_{1}}{\sigma^{2}}\left[1-\left(\tfrac{[\phi(a)-\phi(b)]^{2}}{[1-\Phi(b)+\Phi(a)]^{2}}+\tfrac{\phi(a)\cdot a-\phi(b)\cdot b}{1-\Phi(b)+\Phi(a)}\right)\right]&\text{if }\;\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}\end{cases}$ (14)
    where
    $a=\tfrac{-\sqrt{c_{\chi}-\tilde{y}^{\top}P_{X_{-1}}\tilde{y}}-\beta_{1}\sqrt{X_{1}^{\top}(I_{n}-P_{X_{-1}})X_{1}}}{\sigma},\quad b=\tfrac{\sqrt{c_{\chi}-\tilde{y}^{\top}P_{X_{-1}}\tilde{y}}-\beta_{1}\sqrt{X_{1}^{\top}(I_{n}-P_{X_{-1}})X_{1}}}{\sigma},$ (15)
    and $\phi$ and $\Phi$ denote the density and distribution functions of the standard normal distribution, respectively.
For insight into the two cases in (14), note that if $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$, then the set $R(\tilde{y};X)$ in (13) can more succinctly be characterized as $R(\tilde{y};X):=\left\{Y\in\mathbb{R}^{n}:P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}\right\}$, because $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$ and $P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}$ imply that $Y^{\top}P_{X_{-1}}Y\geq c_{\chi}$ and hence $Y^{\top}P_{X}Y\geq c_{\chi}$. Furthermore, $\tfrac{\partial^{2}}{\partial\beta_{1}^{2}}\ell(\beta_{1};Y\mid P_{X_{-1}}Y=P_{X_{-1}}\tilde{y})=X_{1}^{\top}(I-P_{X_{-1}})X_{1}/\sigma^{2};$ this is simply the (unconditional) Fisher information about $\beta_{1}$ when $\beta_{2},\ldots,\beta_{p}$ are treated as nuisance parameters. Therefore, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}\geq c_{\chi}$, conditioning on the set $R(\tilde{y};X)$ does not affect the Fisher information about $\beta_{1}$. On the other hand, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}$, this more succinct characterization of $R(\tilde{y};X)$ is no longer possible, and $\tfrac{\partial^{2}}{\partial\beta_{1}^{2}}\ell(\beta_{1};Y\mid R(\tilde{y};X))\neq X_{1}^{\top}(I-P_{X_{-1}})X_{1}/\sigma^{2}.$ Thus, when $\tilde{y}^{\top}P_{X_{-1}}\tilde{y}<c_{\chi}$, conditioning on $Y\in R(\tilde{y};X)$ indeed affects the Fisher information about $\beta_{1}$.
[…]
$={\mathbb{E}}\left[\mathcal{I}_{Y|Y\in R(\tilde{Y};X)}(\beta_{1};Y\in R(\tilde{Y};X))\mid\tilde{Y}\in R_{1}(X)\right]\cdot\Pr(\tilde{Y}\in R_{1}(X)).$ (16)
The right-hand side of (2.3) is the product of two terms: the expected leftover Fisher information conditional on rejecting $H_{0}^{1:p}$, and the probability of rejecting $H_{0}^{1:p}$.
To approximate the conditional expectation in (2.3), in the left-hand panel of Fig. 4 we generate 1000 draws of $\tilde{Y}$ according to (1) (with $\beta_{0}=0$) and display the average of (14) over all realizations of $\tilde{Y}$ such that $\tilde{Y}^{\top}P_{X}\tilde{Y}\geq c_{\chi}$. The second term on the right-hand side of (2.3) can be computed explicitly by noting that, for $R_{\rho}(X)$ defined in (12) and $\lambda=\frac{1}{2\sigma^{2}}\beta^{\top}X^{\top}X\beta$, $\Pr(Y\in R_{\rho}(X))=1-F_{p,n\cdot\rho-p-1}(\lambda;F^{-1}_{p,n\cdot\rho-p-1}(0;1-\alpha_{0})),$ where $F_{df_{1},df_{2}}(\lambda;\cdot)$ denotes the cumulative distribution function of a noncentral $F$ random variable with $df_{1}$ and $df_{2}$ degrees of freedom and noncentrality parameter $\lambda$. This quantity is displayed in the middle panel of Fig. 4. Finally, in the right-hand panel of Fig. 4 we display the product of the two terms in the left-hand and center panels, i.e., the right-hand side of (2.3).
For the sample splitting procedure, we plot an analogous set of quantities. Recalling from Proposition 2.3(i) that the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$ is $\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}\right)$, it follows that
${\mathbb{E}}$ $\left[\tfrac{(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}}{\sigma^{2}}\cdot\bm{I}_{R_{\rho}(X^{\mathrm{tr}})}(Y^{\mathrm{tr}})\right]=\tfrac{(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}}{\sigma^{2}}\cdot\Pr(Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})).$ (17)
In the left-hand and center panels of Fig. 4, we display the two terms on the right-hand side of (17), and in the right-hand panel we display their product.
In the left-hand panel of Fig. 4, we see that the selective approach generally results in higher expected leftover Fisher information about $\beta_{1}$, conditional on rejection of $H_{0}^{1:p}$, as compared to sample splitting. Furthermore, the center panel shows that the probability of rejecting $H_{0}^{1:p}$ is always larger using the selective approach than using sample splitting; this is due to the fact that the former uses all $n$ observations whereas the latter uses only $\rho n\leq n$. Finally, the right-hand panel shows that the product of the left and middle panels (representing (2.3) and (17)) is larger for the selective approach than for sample splitting, for all values of $\beta_{1},\ldots,\beta_{p}$ considered.
Remark 2.
We see from (14) that when $\beta_{1}=0$, conditioning on $R(\tilde{Y};X)$ in (13) can increase — and cannot decrease — the leftover Fisher information about $\beta_{1}$ relative to the unconditional Fisher information. This is reflected in the bottom-left panel of Figure 4. However, for larger values of $|\beta_{1}|$, conditioning typically does not increase the leftover Fisher information.
$\beta_{1}=0.5$ $\beta_{1}=0.25$ $\beta_{1}=0$
   $\mathbb{E}$$\big$[Leftover Fisher $\mid\text{rej }\bm{H_{0}^{1:p}}\big]$      Pr(rej $\bm{H_{0}^{1:p}}$)       $\mathbb{E}$$\big$[Leftover Fisher $\mid\text{rej }\bm{H_{0}^{1:p}}\big]\bm{\cdot}$ Pr(rej $\bm{H_{0}^{1:p}}$)
Figure 4: We construct an $n\times p$ orthogonal design matrix $X$ with $n=100$ and $p=5$. Then, for $\beta_{1}\in\{0.5,0.25,0\}$ and for a range of values of $s\in[-1,1]$, we generate 1000 response vectors $\tilde{Y}$ according to (1) with $\sigma=1$, $\beta_{j}=s$ for all $j\in\{2,\ldots,p\}$, and $\beta_{0}=0$. We then compute quantities related to the leftover Fisher information about $\beta_{1}$ for our selective approach, and for sample splitting with split proportions $\rho\in\{0.1,0.5,0.9\}$. Left: For each of the 1000 realizations $\tilde{y}$ of $\tilde{Y}$, we compute the leftover Fisher information $\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))$ for our conditional selective inference procedure and display the average over all datasets such that $\tilde{Y}\in R_{1}(X)$. This approximates ${\mathbb{E}}\left[\mathcal{I}_{Y\mid Y\in R(\tilde{Y};X)}(\beta_{1};Y\in R(\tilde{Y};X))\mid\tilde{Y}\in R_{1}(X)\right]$. Further, for $\rho\in\{0.1,0.5,0.9\}$, we display $(X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}/\sigma^{2}$. Center: We display $\Pr(Y\in R_{1}(X))$, along with $\Pr(Y^{\mathrm{tr}}_{\rho}\in R_{\rho}(X^{\mathrm{tr}}))$ for $\rho\in\{0.1,0.5,0.9\}$. Right: We display the product of the left-hand and center columns, i.e., the right-hand sides of (2.3) and (17).
2.4 A plug-in estimator of the error variance
Proposition 2.1 provides a straightforward way to approximate $p_{H_{0}^{M}\mid E}$} (9) via Monte Carlo, provided that $\sigma^{2}$ is known. In practice, we rarely know the true value of $\sigma^{2}$. The typical solution is to use a plug-in estimator for $\sigma^{2}$; see, for example, Lee et al. (2016, §7). A standard estimator for $\sigma^{2}$ is
$\hat{\sigma}^{2}:=y^{\top}(I_{n-1}-P_{X})y/(n-p-1).$ (18)
Using the notation in Proposition 2.1, this suggests we might estimate (9) as
${p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}}:=\Pr\left(W/Z\geq r\mid W-cZ\geq d/\hat{\sigma}^{2}\right).$ (19)
However, $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} is quite anti-conservative (see the left panel of Fig. 5).
What has gone wrong? Conditional on the event that we reject $H_{0}^{1:p}$ with $F_{H_{0}^{1:p}}$} (3), the estimate $\hat{\sigma}^{2}$ in (18) is biased downward. To understand why, note that there are two reasons we might reject $H_{0}^{1:p}$ using (3): because the numerator is large, or the denominator (which is proportional to $\hat{\sigma}^{2}$) is small. Thus, among datasets for which $H_{0}^{1:p}$ is rejected, $\hat{\sigma}^{2}$ will tend to be small. Furthermore, because (3) and (4) share a denominator, a small value of $\hat{\sigma}^{2}$ is associated with a larger value of (4), leading to a tendency to reject $H_{0}^{M}$ too often. To address this problem, we propose a conditionally “debiased” estimator of the error variance.
Proposition 2.4.
Let $A\sim\chi^{2}_{p}(\lambda)$ and $Z\sim\chi^{2}_{n-p-1}$, where $\lambda=\frac{1}{2\sigma^{2}}\beta^{\top}X^{\top}X\beta$ and $A\perp Z$. Define
${\tilde{\sigma}^{2}}(\lambda):=\frac{Y^{\top}(I_{n-1}-P_{X})Y}{{\mathbb{E}}\left(Z\mid A\geq cZ\right)}.$ (20)
Then ${\mathbb{E}}({\tilde{\sigma}^{2}}(\lambda)|Y\in E_{1})=\sigma^{2}$, i.e., ${\tilde{\sigma}^{2}}(\lambda)$ is unbiased for $\sigma^{2}$ conditional on rejection of $H_{0}^{1:p}$.
An immediate consequence of Proposition 2.4 is that under $H_{0}^{1:p}$ (2), ${\tilde{\sigma}^{2}}:={\tilde{\sigma}^{2}}(0)$ is unbiased for $\sigma^{2}$ conditional on rejection of $H_{0}^{1:p}$. Further, under $H_{0}^{1:p}$, the denominator of $\tilde{\sigma}^{2}$} can be easily approximated with Monte Carlo. This leads to the p-value that we use in practice:
${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}:=\Pr\left(W/Z\geq r\mid W-cZ\geq d/{\tilde{\sigma}^{2}}\right).$ (21)
Away from the overall null, $\tilde{\sigma}^{2}$} is biased upward; however, in our experiments we find that this upward bias is modest, and that its effect on the p-values is minimal (see Appendix C). We see in the left panel of Fig. 5 that $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} yields nominal selective Type 1 error control for values of $\alpha$ that might reasonably be used for hypothesis testing (i.e., $\alpha<0.4).$ In the next proposition we establish that $\tilde{\sigma}^{2}$} is consistent for $\sigma^{2}$, conditional on $Y\in E_{1}$.
Proposition 2.5.
Consider the same setting as described in Proposition 2.2 and Remark 1, but without any requirement on $\beta$. Then conditional on $Y_{n}\in E_{1},$ the estimator ${\tilde{\sigma}^{2}}={\tilde{\sigma}^{2}}(0)$ (20) converges in probability to $\sigma^{2},$ i.e., $\forall\varepsilon>0$, $\lim_{n\to\infty}\Pr\left(\left|{\tilde{\sigma}^{2}}-\sigma^{2}\right|>\varepsilon\mid Y_{n}\in E_{1}\right)=0$.
Consistency of $\tilde{\sigma}^{2}$} for $\sigma^{2}$ conditional on $Y_{n}\in E_{1}$ allows us to establish that ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}(Y_{n})$ controls the selective Type 1 error asymptotically.
Proposition 2.6.
Under $H_{0}^{M}:\beta_{M}=0$ and conditional on $Y_{n}\in E_{1}$, the p-value $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21) converges to the $\text{Unif}(0,1)$ distribution as $n\to\infty$. That is, for all $\alpha^{\prime}\in(0,1)$, $\lim_{n\to\infty}\Pr_{\beta_{M}=0}\left({p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}(Y_{n})\leq\alpha^{\prime}\mid Y_{n}\in E_{1}\right)=\alpha^{\prime}$.
3 Selective p-values in a retrospective analysis
In the applied literature, examples abound of analyses where Step 2 was carried out without accounting for Step 1333Here are some examples of published papers that clearly state that the two-step procedure in Box 1 was applied with a standard test in Step 2: (i) “If the groups tested are different, we report the p-values of post hoc tests but omit the p-value of the omnibus test. If the groups did not test as different, only the P value of the omnibus test is reported” (Kula et al., 2024, page 13); (ii) Table 2 of Wang et al. (2024), where NA’s denote hypotheses that were left untested because the overall test was not rejected; (iii) “Post hoc pairwise comparisons were only performed if the omnibus test was significant” (Tams et al., 2020, caption of Figure 5).. If the original data for such an analysis were available to the reader, then they could reanalyze the data using the methods in Section 2. Unfortunately, the raw data are often not available. In this section, we demostrate that one can use the usual software output of a least squares linear regression (e.g., the output of summary(lm(y~x)) in R) to “correct” Step 2 of Box 1 when a standard test of $H_{0}^{M}$ was used. We refer to this as a retrospective analysis, as it can be applied retrospectively to the results of a published paper, without access to the raw data.
Standard software packages for fitting linear models output the proportion of variance explained ($\mathrm{R}^{2}$) and the residual standard error ($\mathrm{RSE}$), defined as
$\mathrm{R}^{2}$ $=1-\frac{y^{\top}(I_{n-1}-P_{X})y}{y^{\top}y},\;\;\;\mathrm{RSE}=\sqrt{\frac{y^{\top}(I_{n-1}-P_{X})y}{n-p-1}},$ (22)
as well as $F_{H_{0}^{M}}$ (4). These three quantities suffice to compute $p_{H_{0}^{M}\mid E}$} (10).
Proposition 3.1.
Let $F_{H_{0}^{M}}$ be defined as in (4), $c$ defined as in (11), and $\mathrm{R}^{2}$ and $\mathrm{RSE}$ defined as in (22). Define $Z\sim\chi_{n-p-1}^{2}$ and $W\sim\chi_{m}^{2}$, where $Z\perp W$. When $\sigma^{2}$ is known, the p-value in (10) can equivalently be written as
${p_{H_{0}^{M}\mid E}}=\Pr\Bigg(\tfrac{W}{Z}\geq F_{H_{0}^{M}}\cdot\tfrac{m}{n-p-1}\Big|W-cZ\geq\tfrac{\mathrm{RSE}^{2}}{\sigma^{2}}\cdot\left(F_{H_{0}^{M}}\cdot m-\tfrac{\mathrm{R}^{2}}{1-\mathrm{R}^{2}}\cdot(n-p-1)\right)\Bigg).$ (23)
Approximating $\tilde{\sigma}^{2}$} (20) via Monte Carlo only requires knowledge of $n$, $p$, $\alpha_{0}$, and $y^{\top}(I_{n-1}-P_{X})y=\mathrm{RSE}^{2}\cdot(n-p-1)$; therefore, $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21) can also be computed retrospectively.
In contrast, the selective confidence interval and point estimates described in Appendix B involve quantities that are not generally reported in software output, and thus cannot be easily obtained retrospectively.
4 The special case of ANOVA
F-screening is particularly pervasive in the context of ANOVA, where it dates back almost a century (Fisher, 1935; Lindquist, 1940). As a simple example, suppose that we are interested in how plant height, $Y$, varies between treatments by fertilizers $A$, $B$, and $C$. We fit the model $Y=\beta_{A}\cdot\bm{I}_{A}+\beta_{B}\cdot\bm{I}_{B}+\beta_{C}\cdot\bm{I}_{C}+\epsilon,$ where $\bm{I}_{A}$, $\bm{I}_{B}$, and $\bm{I}_{C}$ are $n$-dimensional indicator variables for treatment with fertilizers $A$, $B$, and $C$. In the language of ANOVA, $H_{0}^{1:p}:\beta_{A}=\beta_{B}=\beta_{C}$ is often referred to as the “omnibus test.” If $H_{0}^{1:p}$ is rejected, then one typically conducts post hoc tests (Everitt & Skrondal, 2010) to see how the expected plant height differs across fertilizers; for instance, we may test $H_{0}:\beta_{A}=\beta_{B}$. This post hoc test can be equivalently expressed as a test of $H_{0}:\beta_{B\hskip 1.42262pt\text{vs.}\hskip 1.42262ptA}=0$ in the reparameterized model
$Y$ $=\beta_{A}\cdot(\bm{I}_{A}+\bm{I}_{B})+(\beta_{B}-\beta_{A})\cdot\bm{I}_{B}+\beta_{C}\cdot\bm{I}_{C}+\epsilon$
$=\beta_{A}\cdot(\bm{I}_{A}+\bm{I}_{B})+\beta_{B\hskip 1.42262pt\text{vs.}\hskip 1.42262ptA}\cdot\bm{I}_{B}+\beta_{C}\cdot\bm{I}_{C}+\epsilon.$ (24)
However, as we have seen in Section 2, a post hoc test of $H_{0}:\beta_{B\hskip 1.42262pt\text{vs.}\hskip 1.42262ptA}=0$ that does not account for the fact that $H_{0}^{1:p}$ was rejected will fail to control the Type 1 error.
The machinery developed in Sections 23 for valid inference in Step 2 of F-screening applies directly in the setting of ANOVA. In fact, retrospective analysis, as detailed in Section 3, is even more straightforward because in the context of ANOVA, it is not necessary to have access to $F_{H_{0}^{M}}$, $\mathrm{R}^{2}$, and $\mathrm{RSE}$ to conduct a retrospective analysis: instead, it suffices to know the number of samples within each group, and the mean and standard deviation of the outcome variable within each group. This is made precise in the following remark.
Remark 3.
Consider a one-way ANOVA with $K$ groups, parameterized analogously to (4). Suppose we have access to only the following quantities, for $k=1,\ldots,K$:
  1. 1.
    $n_{k},$ the number of observations in the $k^{\text{th}}$ group,
  2. 2.
    $\hat{\mu}_{k}$, the sample mean of $Y$ among observations in the $k^{\text{th}}$ group, and
  3. 3.
    $\hat{\sigma}_{k}=\sqrt{\frac{1}{n_{k}-1}\sum_{i\in\text{ group }k}(y_{i}-\hat{\mu}_{k})^{2}}$, the sample standard deviation of $Y$ among observations in the $k^{\text{th}}$ group.
Then $\mathrm{R}^{2}$ (22), $\mathrm{RSE}$ (22), $F_{H_{0}^{k_{1}\hskip 0.7113pt\text{vs.}\hskip 0.7113ptk_{2}}}$ (4), and $F_{H_{0}^{1:K}}$ (3) take the form
$\mathrm{R}^{2}=\tfrac{\text{SSB}}{\text{SSB}+\text{SSW}},\quad\mathrm{RSE}=\sqrt{\tfrac{\text{SSW}}{n-K}},\quad F_{H_{0}^{k_{1}\hskip 0.7113pt\text{vs.}\hskip 0.7113ptk_{2}}}=\tfrac{(n-K)(\hat{\mu}_{k_{1}}-\hat{\mu}_{k_{2}})^{2}}{\text{SSW}\left(\frac{1}{n_{k_{1}}}+\frac{1}{n_{k_{2}}}\right)},\quad F_{H_{0}^{1:K}}=\tfrac{\text{SSB}/(K-1)}{\text{SSW}/(n-K)},$
where we define $n=\sum_{k=1}^{K}n_{k}$, $\bar{y}=\frac{1}{n}\sum_{k=1}^{K}\hat{\mu}_{k}n_{k}$, $\text{SSB}=\sum_{k=1}^{K}n_{k}(\hat{\mu}_{k}-\bar{y})^{2}$, and $\text{SSW}=\sum_{k=1}^{K}(n_{k}-1)\hat{\sigma}_{k}^{2}$. Thus $p_{H_{0}^{1:K}}=1-F_{K-1,n-K}(F_{H_{0}^{1:K}})$, and $p_{H_{0}^{M}\mid E}$} can be computed via (23), and further we can approximate $\tilde{\sigma}^{2}$} and compute $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21).
5 Simulation study
5.1 Type 1 error
We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, and $\beta_{1}=\ldots=\beta_{10}=0$, so that $H_{0}^{1:p}$ in (2) holds. We consider testing $H_{0}^{1}:\beta_{1}=0$ if and only if $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$ using $F_{H_{0}^{1:p}}$} (3). For the subset of $100000$ datasets for which $H_{0}^{1:p}$ is rejected, the left-hand panel of Fig. 5 displays (i) $p_{H_{0}^{M}}$ from (5), which does not account for rejection of $H_{0}^{1:p}$; (ii) $p_{H_{0}^{M}\mid E}$} from (10), which accounts for rejection of $H_{0}^{1:p}$ and requires known variance; (iii) $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} from (19), which uses the standard variance estimator and does not control the selective Type 1 error; and (iv) $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} from (21), which uses a conditionally debiased variance estimator. We see that $p_{H_{0}^{M}\mid E}$} controls selective Type 1 error, as does $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} when $\alpha$ is not large (i.e., in the regime of interest). Because $p_{H_{0}^{M}}$ and $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} do not control the selective Type 1 error, we do not consider them in Section 5.2.
Figure 5: Left: For the simulation study detailed in Section 5.1, we display a $\operatorname{Unif}(0,1)$ quantile-quantile plot of p-values for the test of $H_{0}^{1}:\beta_{1}=0$, conditional on the event that $H_{0}^{1:p}$ (2) is rejected. Here, $H_{0}^{1:p}$ (and hence $H_{0}^{1}$) holds, and so a p-value for $H_{0}^{1}:\beta_{1}=0$ that controls the selective Type 1 error should follow a $\operatorname{Unif}(0,1)$ distribution. If the trace lies below the diagonal, the p-value is anti-conservative. For the subset of simulated datasets for which $H_{0}^{1:p}$ is rejected using $F_{H_{0}^{1:p}}$}, we display $p_{H_{0}^{M}}$ (5), $p_{H_{0}^{M}\mid E}$} (10), $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21), and $p^{\mathrm{\hat{\sigma}^{2}}}_{H_{0}^{M}\mid E}$} (19). Middle: For the simulation setting described in Section 5.2, we plot the power to reject $H_{0}^{1:p}$ at three different significance levels ($\alpha_{0}\in\left\{0.05,0.1,0.5\right\}$) with (i) the $F$-statistic based on all of the data, $F_{H_{0}^{1:p}}$ (3), and (ii) the $F$-statistic based on only the training set from sample splitting, $F_{H_{0}^{1:p}}^{\mathrm{train}}$. Right: For the simulation setting described in Section 5.2, we plot the conditional probability of rejecting $H_{0}^{1}$ at significance level 0.05 (marked with a gray dashed line) given that $H_{0}^{1:p}$ was rejected at three levels of significance ($\alpha_{0}\in\{0.05,0.1,0.5\})$. When we reject $H_{0}^{1:p}$ with $F_{H_{0}^{1:p}}$}, we compute the p-values (i) $p_{H_{0}^{M}\mid E}$} (10), and (ii) $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21). When we reject $H_{0}^{1:p}$ with $F_{H_{0}^{1:p}}^{\mathrm{train}},$ we compute $p_{H_{0}^{M}}^{\text{test}}$.
5.2 Power
We compare the power of our proposed approach for F-screening to an approach that relies on sample splitting (Cox, 1975).
First, we formalize the sample splitting procedure. We randomly partition $n$ independent realizations of (1) into two equally-sized training and test sets. We use the training set to test $H_{0}^{1:p}$ using an F-statistic (which we refer to as $F_{H_{0}^{1:p}}^{\mathrm{train}}$), and, if and only if we reject $H_{0}^{1:p}$, we use the test set to test $H_{0}^{M}$ using an F-statistic, which yields the p-value $p_{H_{0}^{M}}^{\text{test}}$. Since the training and test sets are independent, $p_{H_{0}^{M}}^{\text{test}}$ is uniformly distributed under $H_{0}^{M}$, regardless of the value of $F_{H_{0}^{1:p}}^{\mathrm{train}}$.
Now we return to the simulation set-up. We generate data according to (1) with $n=100$, $p=10$, $\sigma^{2}=1$, $X_{ij}\stackrel{{\scriptstyle\mathclap{iid}}}{{\sim}}\mathcal{N}(0,1)$, $\beta_{2}=\ldots=\beta_{10}=0$, and with $\beta_{1}$ ranging from 0 to $0.8$.
First, we test $H_{0}^{1:p}$ (2) at $\alpha_{0}\in\left\{0.05,0.1,0.5\right\}$ using either (i) all the data, or (ii) just the training data. The middle panel of Fig. 5 displays the proportion of simulated datasets for which $H_{0}^{1:p}$ is rejected. Unsurprisingly, we have substantially more power to reject $H_{0}^{1:p}$ using all the data versus using only the training data.
Next, for the subset of simulated datasets for which we rejected $H_{0}^{1:p}$ in the middle panel of Fig. 5, we test the hypothesis $H_{0}^{1}:\beta_{1}=0$. Simulated datasets for which $H_{0}^{1:p}$ was rejected using all the data are tested using $p_{H_{0}^{M}\mid E}$} (10) and $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21), and those for which $H_{0}^{1:p}$ was rejected using the training data are tested using $p_{H_{0}^{M}}^{\text{test}}$. The right-hand panel of Fig. 5 displays the power to reject $H_{0}^{1}:\beta_{1}=0$, conditional on having rejected $H_{0}^{1:p}$. It is highest for $p_{H_{0}^{M}\mid E}$} (10), slightly lower for $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21), and substantially lower for $p_{H_{0}^{M}}^{\text{test}}$. The power curves for $p_{H_{0}^{M}}^{\text{test}}$ for the three values of $\alpha_{0}$ are nearly identical, since the training data used to test $H_{0}^{1:p}$ is independent of the test data used to test $H_{0}^{1}$, and $\alpha_{0}$ appears only in the test of $H_{0}^{1:p}$. By contrast, the selective p-values $p_{H_{0}^{M}\mid E}$} and $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} shift substantially according to the value of $\alpha_{0}$: when $\alpha_{0}$ is small, the conditional probability of rejecting $H_{0}^{1}:\beta_{1}=0$ is smaller, since relatively little signal for $\beta_{1}$ is available for inference after conditioning.
It is worth noting that the benefits of our selective p-value over the sample splitting p-value extend beyond the empirical results seen in Fig. 5. Depending on how sample splitting is performed, the testable hypotheses may differ between train and test sets (e.g., if rank deficiency emerges from random partitioning). Also, it may be hard in an applied setting to convince a collaborator to use only part of their data to test $H_{0}^{1:p}$ and to reserve the remainder for testing $H_{0}^{M}$. By contrast, our selective p-values enable the collaborator to use all of the data to test $H_{0}^{1:p}$, thereby yielding a result for the test of $H_{0}^{1:p}$ that coincides with the standard (uncorrected) F-screening procedure. Furthermore, the sample splitting p-values cannot be computed retrospectively using summary statistics, as in Section 3.
5.3 Confidence intervals
We now investigate coverage and width of the confidence intervals derived in Appendix B.
To investigate confidence interval coverage, we generate data according to (1) with $\beta_{1}=\ldots=\beta_{p}=0$, $n=100$, $p=5$, and $\sigma^{2}=1$ until we obtain $1000$ datasets $(X,Y)$ for which we reject $H_{0}^{1:p}$ (2) at level $\alpha_{0}=0.05$. Then, for a fixed confidence level (defined as $(1-\alpha)$ for $\alpha\in(0,1)$), we compute, for each of the $1000$ datasets: (i) the standard confidence interval for $\beta_{1}$ that does not account for selection, (ii) the selective confidence interval for $\beta_{1}$ using the true value of $\sigma^{2}$ (B.13), and (iii) the selective confidence interval for $\beta_{1}$ using a conditionally debiased estimate of $\sigma^{2}$ (B.16).
As shown in the left panel of Fig. 6, the standard confidence intervals for $\beta_{1}$ that do not account for rejection of $H_{0}^{1:p}$ (2) exhibit severe under-coverage for all confidence levels. In contrast, the selective confidence interval that utilizes the true value of $\sigma^{2}$ achieves exact nominal coverage across all confidence levels, while the selective confidence interval that employs the debiased estimate of ${\sigma}^{2}$ attains nominal coverage for confidence levels corresponding to $(1-\alpha)\geq 0.5$ but slightly undercovers for $(1-\alpha)<0.5$. We note that, in practice, $(1-\alpha)$ is typically chosen to be much larger than 0.5.
Figure 6: For the setting described in Section 5.3, we construct confidence intervals for $\beta_{1}$ in (1) according to Appendix B, conditional on rejection of $H_{0}^{1:p}$ (2) with $F_{H_{0}^{1:p}}$} (3). Left: We plot the coverage of the selective and standard confidence intervals for the datasets for which $H_{0}^{1:p}$ (2) is rejected, where the significance level $\alpha$ varies between 0 and 1, and $\beta_{1}=\ldots=\beta_{p}=0$. Since the standard intervals do not achieve the nominal coverage, we do not display them in the middle and right panels. Middle: We plot the mean width of the selective confidence intervals for the datasets for which $H_{0}^{1:p}$ (2) is rejected, where $\beta_{1}$ varies and $n=100$. Right: We plot the mean width of the selective confidence intervals for which $H_{0}^{1:p}$ (2) is rejected for $\beta_{1}=0$ as $n$ varies.
To examine confidence interval width, we first vary $\beta_{1}\in\left[-1,1\right]$ while holding $n=100$ constant. We set $p=5$, $\sigma^{2}=1$, and $\beta_{2}=\dots=\beta_{5}=0$. For each fixed value of $\beta_{1}\in\left[-1,1\right]$, we generate data according to (1) until we obtain 1000 datasets where the overall hypothesis test is rejected. For each dataset, we compute (i) the selective confidence interval using the true value of $\sigma^{2}$, and (ii) the selective confidence interval using the debiased $\tilde{\sigma}^{2}$. (We do not depict the standard confidence interval, as we showed in the left-hand panel of Fig. 6 that it is invalid.) Finally, we compute the average confidence interval width across the 1000 datasets and plot these values as a function of $\beta_{1}$ in the middle panel of Fig. 6. In the right-hand panel, we repeat the analysis in the middle panel, but with $\beta_{1}=0$ and $n\in\{50,\ldots,500\}.$
We now comment on the shape of the curves shown in the middle panel of Fig. 6. The selective confidence intervals—whether computed using the true value of $\sigma^{2}$ or the debiased estimate of $\sigma^{2}$—exhibit a striking shape: they are narrow both when $\beta_{1}\approx 0$ and when $|\beta_{1}|$ is large. This pattern can be understood by separately considering the cases where (i) $\left\lvert\beta_{1}\right\rvert$ is large, (ii) $\left\lvert\beta_{1}\right\rvert$ is moderate, and (iii) $\left\lvert\beta_{1}\right\rvert$ is small. When $\left\lvert\beta_{1}\right\rvert$ is large, the event $E_{1}$ (6) is highly likely, in which case conditional inference roughly coincides with standard inference, and so the intervals are not appreciably wider than the standard (non-selective) intervals. When $\left\lvert\beta_{1}\right\rvert$ is moderate, the event $E_{1}$ (6) is only moderately likely, and when it occurs, it often happens that the data falls near the boundary of the selection region, and little signal regarding $\beta_{1}$ remains for inference, leading to wide confidence intervals. Finally, when $\left\lvert\beta_{1}\right\rvert$ is small, the event $E_{1}$ (6) is unlikely to occur; however, when it does occur, selection is often driven by (chance) association between $X_{2},\ldots,X_{p}$ and $Y$, in which case sufficient signal about $\beta_{1}$ remains for inference, leading to moderately narrow confidence intervals.
6 Data analysis
In this section, we demonstrate a valid data analysis involving F-screening, both with and without access to the full data. First we re-analyze the data of Xu & Wu (2018), and report selective p-values, confidence intervals, and point estimates for each coefficient. Then we conduct a retrospective analysis of Keil et al. (2023) and compute the selective p-values using only the standard ANOVA outputs because we do not have access to their data. In both analyses, we compare the selective quantities to their standard counterparts.
6.1 Prospective inference
Xu & Wu (2018) used the publicly-available National Health and Nutrition Examination Survey (NHANES) dataset from Centers for Disease Control and Prevention (2024) (CDC) to investigate possible changes in femoral neck bone mineral density (BMD)  [^4] BMD is coded as “DXXNKBMD” in the NHANES dataset. across the years 2005 to 2014 in US adults over the age of 30. In the NHANES dataset, years are binned into 2005–2006, 2007–2008, 2009–2010, 2013–2014; no data from 2011–2012 is available for BMD. We restrict our analysis to the $n=7135$ males in the dataset, and consider the model
$\text{BMD}=\beta_{\text{AGE}}\cdot\text{AGE}+\beta_{\text{5--6}}\cdot\bm{I}_{\text{5--6}}+\beta_{\text{7--8}}\cdot\bm{I}_{\text{7--8}}+\beta_{\text{9--10}}\cdot\bm{I}_{\text{9--10}}+\beta_{\text{13--14}}\cdot\bm{I}_{\text{13--14}}+\epsilon,$ (25)
where $\bm{I}_{\text{5--6}}$, $\bm{I}_{\text{7--8}}$, $\bm{I}_{\text{9--10}}$, and $\bm{I}_{\text{13--14}}$ are indicator variables for the time periods 2005–2006, 2007–2008, 2009–2010, and 2013–2014, respectively. Note that (25) contains age, as reported in the NHANES dataset, as a covariate. For convenience, we regress out age: let $U\in\mathbb{R}^{7135\times 7134}$ such that $UU^{\top}:=I_{7135}-P_{\text{AGE}}$ and $U^{\top}U=I_{7135}$, and consider the model
$U^{\top}\text{BMD}=\beta_{\text{5--6}}\cdot U^{\top}\bm{I}_{\text{5--6}}+\beta_{\text{7--8}}\cdot U^{\top}\bm{I}_{\text{7--8}}+\beta_{\text{9--10}}\cdot U^{\top}\bm{I}_{\text{9--10}}+\beta_{\text{13--14}}\cdot U^{\top}\bm{I}_{\text{13--14}}+U^{\top}\epsilon.$ (26)
In this setting, the overall null hypothesis is
$H_{0}^{1:p}:\beta_{\text{5--6}}=\beta_{\text{7--8}}=\beta_{\text{9--10}}=\beta_{\text{13--14}}=0.$ (27)
In line with Xu & Wu (2018), we test whether the mean age-adjusted BMD differs between time bins (05–06 versus 07–08, 05–06 versus 09–10, 05–06 versus 13–14, 07–08 versus 09–10, 07–08 versus 13–14, 09–10 versus 13–14). To compare mean age-adjusted BMD in 2005–2006 to mean age-adjusted BMD in 2007–2008, we reparametrize (26) as
$U^{\top}\text{BMD}$ $=\beta_{\text{5--6}}\cdot\left(U^{\top}\bm{I}_{\text{5--6}}+U^{\top}\bm{I}_{\text{7--8}}\right)$
$\hskip 14.22636pt+\beta_{\text{7--8 vs. 5--6}}\cdot U^{\top}\bm{I}_{\text{7--8}}+\beta_{\text{9--10}}\cdot U^{\top}\bm{I}_{\text{9--10}}+\beta_{\text{13--14}}\cdot U^{\top}\bm{I}_{\text{13--14}}+U^{\top}\epsilon,$ (28)
and test $H_{0}^{\text{ 7--8 vs. 5--6}}:\beta_{\text{7--8 vs. 5--6}}=0.$ The other five pairwise tests are conducted similarly. Results are shown in Table 1. In this case, accounting for the rejection of $H_{0}^{1:p}$ (27) has virtually no impact—the selective and standard quantities are identical up three decimals.
Standard Selective
Years Estimate 95% CI p-value Estimate 95% CI p-value
05–06 vs. 07–08 -0.004 (-0.013, 0.005) 0.423 -0.004 (-0.013, 0.005) 0.423
05–06 vs. 09–10 -0.003 (-0.012, 0.006) 0.473 -0.003 (-0.012, 0.006) 0.473
05–06 vs. 13–14 -0.020 (-0.029, -0.010) 0.000 -0.020 (-0.029, -0.010) 0.000
07–08 vs. 09–10 0.000 (-0.008, 0.009) 0.923 0.000 (-0.008, 0.009) 0.923
07–08 vs. 13–14 -0.016 (-0.025, -0.007) 0.000 -0.016 (-0.025, -0.007) 0.000
09–10 vs. 13–14 -0.017 (-0.025, -0.008) 0.000 -0.017 -(0.025, -0.008) 0.000
Table 1: For NHANES BMD data, we perform F-screening for the model in (25), and compare selective point estimates, confidence intervals, and p-values with standard counterparts following the rejection of the overall $F$-test of (27). The statistics in the “Standard” column do not account for rejection of (27) (see Box 1), whereas statistics in the “Selective” column do account for it.
6.2 Retrospective inference
Keil et al. (2023) explores the relationship between sleep and cognitive impairment in older adults. The authors considered four response variables: (i) Center for Epidemiologic Studies-Depression Scale (CES-D), (ii) sleep duration, (iii) Mini-Mental State Examination (MMSE), and (iv) Mattis Dementia Rating Scale (MDRS) (see references in Keil et al. (2023) for details). For each response variable corresponding to the rows in Fig. 1, Keil et al. (2023) fit a model of the form $Y=\beta_{1}\cdot\bm{I}_{<65}+\beta_{2}\cdot\bm{I}_{\text{65--85}}+\beta_{3}\cdot\bm{I}_{>85}+\epsilon,$ where $\bm{I}_{<65}$, $\bm{I}_{\text{65--85}}$, and $\bm{I}_{>85}$ denote indicator variables for the age group (in years) of $n=826$ study participants. The second-to-last column of Fig. 1 displays the p-values for a test of $H_{0}^{1:p}:\beta_{1}=\beta_{2}=\beta_{3}=0$, and the last column reports p-values for standard tests of $H_{0}^{j}:\beta_{j}=0$ for $j=1,2,3$.
Keil et al. (2023) are not explicit about whether they would have tested $H_{0}^{j}$, $j=1,2,3$ had they not rejected $H_{0}^{1:p}$, and it is not possible to ascertain their intentions from their results, since in their Table 1 (a subset of which is displayed in our Fig. 1), all hypotheses of the form $H_{0}^{1:p}$ tested with ANOVA are rejected. Notably, they refer to tests of $H_{0}^{j}$ as “post hoc tests”; this phrasing is typically reserved in the applied literature to refer to tests that are conducted after rejecting $H_{0}^{1:p}$ (Everitt & Skrondal, 2010). Thus, for the sake of demonstration we shall assume that they would not have tested $H_{0}^{j}$ for $j=1,2,3$ in their Table 1 if $H_{0}^{1:p}$ had not been rejected.
Accordingly, we apply the methods in Section 3 and 4 to conduct a retrospective analysis, using only the summary statistics reported in Fig. 1. For each of the four outcomes in Fig. 1, we compute $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21) for $H_{0}^{j}:\beta_{j}=0$, for $j=1,2,3$, alongside the standard p-value in (5) which does not account for the fact that $H_{0}^{1:p}$ was rejected. The standard p-values that we compute match those in Fig. 1 if the Šidák correction is applied (Šidák, 1967).
The results are displayed in Table 2. Generally, the values of $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21) are equal to or larger than the corresponding values of $p_{H_{0}^{M}}$ (5). For example, the p-value for the mean difference in CES-D between the under-65 age group and the 65-85 age group is 0.003 using $p_{H_{0}^{M}}$ versus 0.117 using $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}, and the p-value for the mean difference in CES-D between the 65-85 age group and over-85 age group is 0.016 versus 0.099. At the significance level of 0.05 used by Keil et al. (2023), the selective p-value $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} does not lead to rejection for those two post hoc tests, whereas the standard p-value $p_{H_{0}^{M}}$ does.
        CES-D     Sleep     MMSE     MDRS
        $p_{H_{0}^{M}}$ $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}     $p_{H_{0}^{M}}$ $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}     $p_{H_{0}^{M}}$ $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}     $p_{H_{0}^{M}}$ $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$}
    <65<65 vs. 65-85     0.003 0.117     0.001 0.001     0.000 0.000     0.003 0.003
    <65<65 vs. $\geq 85$     0.397 0.397     0.000 0.003     0.000 0.000     0.000 0.000
    65-85 vs. $\geq 85$     0.016 0.099     0.060 0.061     0.000 0.000     0.000 0.000
Table 2: For the retrospective analysis conducted on the linear regression outputs from Keil et al. (2023) (Fig. 1), we display the selective ($p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21)) and standard ($p_{H_{0}^{M}}$ in (5)) p-values from four analyses, each involving a separate outcome variable (CES-D, Sleep, MMSE, MDRS). For each of these outcome variables, we replicate the one-way ANOVA and post hoc tests conducted in Keil et al. (2023), which are reported in the last two columns of the table in Fig. 1. For the post hoc tests, we report $p_{H_{0}^{M}}$ (5) and $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} (21). If we were to apply the Šidák correction (Šidák, 1967) to the values of $p_{H_{0}^{M}}$ (5) in this table, they would match the p-values in the last column of Fig. 1.
7 Discussion
We have shown that a common practice in applied statistics of reporting the outputs of a linear regression model if and only if an overall $F$-test is rejected leads to invalid inference. We have proposed an easily-implemented approach, based on conditional selective inference, that remedies the problem. We present a toolbox for valid inference in this setting that contains the most critical linear model outputs: p-values, confidence intervals, and point estimates. Furthermore, our selective p-values can be computed even without access to the raw data, using only the standard outputs of a multiple linear regression.
Some natural questions that we leave for future work are as follows:
  • Our results relied on an assumption that the errors are Gaussian. Can this be relaxed using ideas from e.g. Tian & Taylor (2017); Panigrahi (2023)? (We note, of course, that standard inference for a linear model also assumes Gaussianity in finite samples—thus, reliance on Gaussianity is only a very minor limitation of our method.)
  • Can these results be extended beyond least squares linear regression, e.g. to generalized linear models, using ideas from Tian & Taylor (2018)?
  • How should our selective p-values be adjusted to account for multiple comparisons? For instance, how can $p_{H_{0}^{M}\mid E}$} for $M\in\{M_{1},\ldots,M_{L}\}$ be corrected in order to control family-wise selective Type 1 error rate? An extension of Scheffé’s correction (Scheffé, 1953) to the selective setting may be possible.
An R package called lmFScreen, implementing the methods proposed in this paper, is available at https://github.com/mcgougho/lmFScreen. R scripts to reproduce all numerical results can be found at https://github.com/mcgougho/lmFScreen-paper.
Acknowledgements
The authors gratefully acknowledge funding from NSF DMS 2322920, NSF DMS 2514344, NIH 5P30DA048736, and ONR N00014-23-1-2589 to DW; NSF DMS-2303371, the Pacific Institute for the Mathematical Sciences, the Simons Foundation, and the eScience Institute at the University of Washington to DK; and NSF Graduate Research Fellowship Grant No. DGE-2140004 to OM.
Appendix A Connections to multiple testing
A.1 Family-wise error rate control does not imply selective Type 1 error rate control
One might wonder whether the problem addressed in this paper could be more simply tackled by controlling the family-wise error rate (FWER). FWER control has been studied for almost a century (Tukey, 1949; Scheffé, 1953; Dunn, 1961; Holm, 1979; Marcus et al., 1976; Hochberg & Tamhane, 1987), and continues to be an area of active research: for instance, the “PoSI” proposal of Berk et al. (2013)—foundational to the field of selective inference—targets the FWER.
However, we see in the right-hand panel of Fig. A.1 that, in the context of F-screening in a setting with $p=3$, control of the FWER does not imply control of the selective Type 1 error rate. In particular, though Bonferroni, Scheffé, and closed testing corrections are guaranteed to control FWER, they do not control the Type 1 error rate conditional on rejection of $H_{0}^{1:3}$.
Figure A.1: We simulate data according to (1) with $n=100$, $p=3$, $\sigma^{2}=1$, and $\beta_{1}=\beta_{2}=\beta_{3}=0$, so that $H_{0}^{1:p}$ (2) holds. For the test of $H_{0}^{1}:\beta_{1}=0$, we report (i) the uncorrected p-value that does not account for F-screening, i.e. $p_{H_{0}^{M}}$ in (5); (ii) $2p_{H_{0}^{M}}$, resulting from a Bonferroni correction to $p_{H_{0}^{M}}$ that accounts for the fact that two tests ($H_{0}^{1:p}$ and $H_{0}^{1}$) were conducted; (iii) $3p_{H_{0}^{M}}$, a Bonferroni correction that accounts for the fact that three post hoc tests ($H_{0}^{j}:\beta_{j}=0$ for $j\in\left\{1,2,3\right\}$) could have been conducted; (iv) the p-value that results from using Scheffé’s method; (v) the p-value associated with the closed testing procedure as described in Section A.2; and (vi) the selective p-value (10) that accounts for F-screening and uses the true value of error variance. Left: For each of 100,000 simulated datasets, the $Q$-plot displays the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ follows a $\operatorname{Unif}(0,1)$ distribution, while the selective p-values are extremely conservative because all results for all simulated datasets are displayed, regardless of rejection of $H_{0}^{1:p}$. Since we are only considering a single test, the Bonferroni-corrected, Scheffé-corrected, and closed-testing p-values are also conservative. Right: For the 5% of the 100,000 simulated datasets for which $H_{0}^{1:p}$ is rejected at level $\alpha_{0}=0.05$, we display the p-values for $H_{0}^{1}$. The uncorrected p-value $p_{H_{0}^{M}}$ and the closed testing p-value are anti-conservative, as are the Bonferroni-corrected and Scheffé-corrected p-values when small quantiles are considered. However, the selective p-value follows a $\operatorname{Unif}(0,1)$ distribution.
Why is this the case? Consider the setting of Fig. A.1. There are three possibilities associated with the null hypotheses $H_{0}^{1:3}:\beta_{1:3}=0$ and $H_{0}^{1}:\beta_{1}=0$: (i) both are true; (ii) $H_{0}^{1:3}$ is false, but $H_{0}^{1}$ is true; or (iii) both are false. In (iii), a Type 1 error is impossible, so we consider only (i) and (ii). Let $\delta_{1:3}$ be the decision function associated with the test of $H_{0}^{1:3}$: it is a function of the (random) data and takes value $1$ when we reject $H_{0}^{1:3}$ and value 0 otherwise; let $\delta_{1}$ similarly denote the decision function associated with testing $H_{0}^{1}$. In scenario (i), the FWER is ${\mathbb{E}}_{H_{0}^{1:3}}\left[\max\{\delta_{1:3},\delta_{1}\}\right]$. In scenario (ii), the FWER is ${\mathbb{E}}_{H_{0}^{1}}\left[\delta_{1}\right]$. By contrast, in scenario (i) the selective Type 1 error — which is controlled by our procedure — is ${\mathbb{E}}_{H_{0}^{1:3}}\left[\delta_{1}\mid\delta_{1:3}=1\right]$, while in scenario (ii) it is ${\mathbb{E}}_{H_{0}^{1}}\left[\delta_{1}\mid\delta_{1:3}=1\right]$. Clearly, the former targets are not the same as the latter targets, and control of one does not imply control of the other.
A.2 Connection to closed testing
At first glance, there appears to be a connection between the F-screening procedure described in Box 1, and so-called “closed testing procedures” (Marcus et al., 1976).
We begin by describing the closed testing procedure of Marcus et al. (1976), which controls the FWER. Let $H_{1},H_{2},\dots,H_{q}$ denote the set of elementary hypotheses to be tested. In addition to the elementary hypotheses, we also consider all possible intersections $H_{I}=\cap_{i\in I}H_{i}$ for all nonempty $I\subseteq\left\{1,2,\dots,q\right\}$. Next, we record the results of a (preliminary) test at level $\alpha$ for each of the hypotheses (including the elementary and intersection hypotheses). We then consider an elementary hypothesis $H_{i}$ to be rejected if and only if the preliminary test $H_{I}$ is rejected for all $H_{I}$ where $i\in I$.
How can we apply closed testing in our setting? First, consider elementary hypotheses $H_{0}^{M}:\beta_{M}=0$ and $H_{0}^{-M}:\beta_{-M}=0$. In order to conduct a closed testing procedure, the only additional hypothesis we need to consider is $H_{0}^{1:p}=H_{0}^{M}\cap H_{0}^{-M}$. In order to reject $H_{0}^{M}$, closed testing dictates that our preliminary tests must have rejected $H_{0}^{1:p}$ and $H_{0}^{M}$; of course, we can conduct the preliminary test of the former before we conduct the preliminary test of the latter. Thus, in this setting, closed testing bears a resemblance to the procedure in Box 1. The p-value associated with the test of $H_{0}^{M}$ in this particular instance of closed-testing is the maximum of the two p-values associated with the preliminary tests of $H_{0}^{1:p}$ and $H_{0}^{M}$. In Fig. A.1, we plot this p-value in the case that $p=3$ and the test of interest in Step 2 of Box 1 is $H_{0}^{1}:\beta_{1}=0$. We see that the closed testing p-value does not control the selective Type 1 error.
How can we reconcile the FWER guarantees of closed testing with the empirical results depicted in the right-hand panel of Figure A.1? The issue hinges on how we “count” the decision regarding $H_{0}^{M}$ in scenarios wherein we fail to reject $H_{0}^{1:p}$. We illustrate this point with a simple example, in which $H_{0}^{1:p}$ holds. Suppose that one hundred researchers independently collect data and then perform closed testing at level $\alpha$. This entails conducting preliminary tests of $H_{0}^{1:p}$, $H_{0}^{M}$, and $H_{0}^{-M}$. Suppose that five of the researchers’ preliminary tests reject $H_{0}^{1:p}$; of these five, four of their preliminary tests also reject $H_{0}^{M}$. Then, the closed testing procedure dictates that these four researchers should reject $H_{0}^{M}$. In this setting, the family-wise error proportion is $5/100=5\%$, because five of the 100 researchers falsely rejected at least one null hypothesis: this is consistent with the fact that closed testing guarantees FWER control. In contrast, the resulting selective Type 1 error proportion — defined as the proportion of researchers who rejected $H_{0}^{M}$ among those who rejected $H_{0}^{1:p}$ — is $4/5=80\%$; this is because closed testing does not control the selective Type 1 error rate.
While other work on closed testing has built upon the proposal of Marcus et al. (1976) and has considered other targets besides FWER (Goeman & Solari, 2011), this literature does not target the selective Type 1 error rate that is the focus of our paper.
Appendix B Selective confidence intervals and point estimates
In this appendix, we derive confidence intervals and point estimates for the regression coefficients that are valid conditional on rejecting $H_{0}^{1:p}$ (2).
B.1 Point estimates
In this section we will develop an approach to find the conditional maximum likelihood estimate (MLE) (conditional on rejecting $H_{0}^{1:p}$) of $\beta_{j}$ under (1). First, we will arrive at a convenient expression for the conditional density. Once that is obtained, we will maximize the associated conditional likelihood with respect to $\beta_{j}$.
Unfortunately, we encounter an immediate challenge in finding the conditional density of $Y$ with respect to $\beta_{j}$ given $Y\in E_{1}$: the probability of the event $E_{1}$ depends on nuisance parameters $\beta_{-j}$ through the projection $P_{X_{-j}}Y$. To eliminate this dependence, we additionally condition on the event $\tilde{E}(\tilde{y})=\{Y:P_{X_{-j}}Y=P_{X_{-j}}\tilde{y}\}$ (where $\tilde{y}$ is a realization of $Y$); this fixes the component of $Y$ that lies in the column space of $X_{-j}$.
Proposition B.1.
Let $Y\sim\mathcal{N}_{n-1}(X\beta,\sigma^{2}I_{n-1})$, and define
$D=\begin{bmatrix}I_{n-1}-P_{X}\\ P_{X}-P_{X_{-j}}\\ P_{X_{-j}}\end{bmatrix}=\begin{bmatrix}D_{1}\\ D_{2}\\ D_{3}\end{bmatrix}$.
Define new random variables $\tilde{Z}=D_{1}Y$, $\tilde{W}=D_{2}Y$, and $\tilde{V}=D_{3}Y$ (note that $\tilde{Z}^{\top}\tilde{Z}/\sigma^{2}\overset{d}{=}Z$ and $\tilde{W}^{\top}\tilde{W}/\sigma^{2}\overset{d}{=}W$, for $Z$ and $W$ that appear in Proposition 2.1). Let $c$ be as defined in (11) and $E_{1}$ as defined in (6), let $\tilde{y}$ denote a realization of $Y$, and define $\tilde{d}=P_{X_{-j}}\tilde{y}$. Define $\tilde{E}(\tilde{y})=\left\{Y:P_{X_{-j}}Y=P_{X_{-j}}\tilde{y}\right\}$.
[…]
$f_{Y\mid Y\in E_{1}\cap\tilde{E}(\tilde{y})}(y)=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{y^{\top}D_{2}y-y^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq\tilde{d}^{\top}\tilde{d}\right).}$, (B.1)
where the marginal densities of $\tilde{Z}$ and $\tilde{W}$ are given in (B.8) and $\bm{I}_{y}\{\cdot\}$ denotes the indicator that $y$ satifies the condition inside the brackets.
Proof.
[…]
$Y\in E_{1}$ $\iff\frac{Y^{\top}P_{X}Y}{Y^{\top}\left(I_{n-1}-P_{X}\right)Y}\geq c$ (B.2)
$\iff\frac{Y^{\top}(P_{X}-P_{X_{-j}})Y+Y^{\top}P_{X_{-j}}Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq c$
$\iff\frac{\tilde{W}^{\top}\tilde{W}+\tilde{V}^{\top}\tilde{V}}{\tilde{Z}^{\top}\tilde{Z}}\geq c$
$\iff\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{V}^{\top}\tilde{V}$,
[…]
$\iff P_{X_{-j}}Y=\tilde{d}$
$\iff\tilde{V}=\tilde{d}$.
[…]
$Y\in E_{1}\cap\tilde{E}(\tilde{y})\iff\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\text{ and }\tilde{V}=\tilde{d}$. (B.4)
[…]
$f_{Y}(y)=f_{\tilde{Z},\tilde{W},\tilde{V}}\left(Dy\right).$ (B.5)
[…]
$f_{Y\mid Y\in\tilde{E}(\tilde{y})}(y)$ $=f_{Y\mid\tilde{V}=\tilde{d}}(y)$ (B.6)
$=f_{\tilde{Z},\tilde{W},\tilde{V}\mid\tilde{V}=\tilde{d}}(Dy)$
$=f_{\tilde{Z}\mid\tilde{V}=\tilde{d}}\left(D_{1}y\right)f_{\tilde{W}\mid\tilde{V}=\tilde{d}}\left(D_{2}y\right)f_{\tilde{V}\mid\tilde{V}=\tilde{d}}\left(D_{3}y\right)$
$=f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)f_{\tilde{V}\mid\tilde{V}=\tilde{d}}\left(D_{3}y\right)$
$=f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y),$
where the first equality follows from (B.3), the second equality follows from (B.5), the third from mutual independence of $\tilde{Z}$, $\tilde{W}$, and $\tilde{V}$ per Lemma E.1, the fourth from the independence of $\tilde{Z}$ from $\tilde{V}$ and of $\tilde{W}$ from $\tilde{V}$, and the last equality follows because the conditional distribution of $\tilde{V}\mid\tilde{V}=\tilde{d}$ is simply an atom at $\tilde{V}=\tilde{d}$.
[…]
$f_{Y\mid Y\in E_{1}\cap\tilde{E}(\tilde{y})}(y)$ $=f_{Y\mid\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d},\tilde{V}=\tilde{d}}(y)$ (B.7)
$=f_{\left\{Y\mid\tilde{V}=\tilde{d}\right\}\mid\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}}(y)$
$=\frac{f_{Y\mid\tilde{V}=\tilde{d}}(y)\bm{I}_{\left\{y\;:\;y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)}$
$=\frac{f_{\tilde{Z}}\left(D_{1}y\right)f_{\tilde{W}}\left(D_{2}y\right)\bm{I}_{\left\{D_{3}y=\tilde{d}\right\}}(y)\bm{I}_{\left\{y\;:\;y^{\top}D_{2}y-cy^{\top}D_{1}y\geq-\tilde{d}^{\top}\tilde{d}\right\}}(y)}{\Pr\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-\tilde{d}^{\top}\tilde{d}\right)}$,
where the first equality follows from (B.4), the second equality follows because sequential conditioning is equivalent to simultaneous conditioning, the third equality from the standard definition of conditional density and the fact that $\tilde{W}=D_{2}Y$ and $\tilde{Z}=D_{1}Y$ from the statement of the proposition, and the fourth by substituting the result of (B.6). ∎
Now, we wish to view (B.7) as a likelihood, i.e., a function of $\beta_{j}$ with the realization of the data $y$ fixed. First, note that because $Y\sim\mathcal{N}_{n-1}\left(X\beta,\sigma^{2}I_{n-1}\right)$,
$\tilde{Z}$ $=\left(I_{n-1}-P_{X}\right)Y\sim\mathcal{N}_{n-1}\left(0_{n-1},\sigma^{2}\left(I_{n-1}-P_{X}\right)\right),$ (B.8)
$\tilde{W}$ $=\left(P_{X}-P_{X_{-j}}\right)Y\sim\mathcal{N}_{n-1}\left(\beta_{j}\left(P_{X}-P_{X_{-j}}\right)X_{j},\sigma^{2}\left(P_{X}-P_{X_{-j}}\right)\right).$
By inspection, we can observe that the distribution of $\tilde{Z}$ is invariant to $\beta_{j}$, while $\tilde{W}$ involves $\beta_{j}$ but none of $\beta_{-j}$. When (B.7) is viewed as a conditional likelihood under the data realization $y$, the indicators must be satisfied (since we observed the data), and maximizing it in $\beta_{j}$ is equivalent to maximizing
$L(\beta_{j}):=\frac{f_{\tilde{W}}(\beta_{j};\tilde{w}(y))}{\Pr_{\beta_{j}}\left(\tilde{W}^{\top}\tilde{W}-c\tilde{Z}^{\top}\tilde{Z}\geq-y^{\top}P_{X_{-j}}y\right)},$ (B.9)
where $\tilde{w}(y)=(P_{X}-P_{X_{-j}})y$, and we write $\Pr_{\beta_{j}}$ to emphasize that the probability depends upon $\beta_{j}$.
Recall that we observe $y$ and hence can compute $\tilde{w}(y)$ and $y^{\top}P_{X_{-j}}y.$ Assuming $\sigma^{2}$ is known, then for any candidate value $\beta_{j}$, the numerator can be computed exactly, and the denominator can be approximated via Monte Carlo. Therefore, to approximate the conditional MLE, we numerically maximize the logarithm of the ratio in (B.9) with respect to $\beta_{j}$. In practice, we will plug in the estimate $\tilde{\sigma}^{2}$} (20) in order to maximize (B.9) with respect to $\beta_{j}$.
We now comment on Fig. B.1. When the overall null $H_{0}^{1:p}$ (2) holds and $H_{0}^{1:p}$ is rejected, the standard MLE has a much larger spread than the conditional MLE derived in this section, and indeed the mode of the former is not even at 0. This is because the overall null is rejected when the F-statistic is sufficiently large, which tends to occur when $\hat{\beta}_{1}$ and/or $\hat{\beta}_{2}$ is large. Thus, the standard MLE for $\beta_{1}$ will tend to be far from zero conditional on the rejection of the overall null. The selective estimate, on the other hand, adjusts for the selection event, and therefore has its mode at the true value 0 and is more tightly concentrated around it.
Figure B.1: We generate datasets $(X,Y)$ with $n=10$, $p=2$, and $\sigma^{2}=2$ under the overall null $H_{0}^{1:p}$ until we have $1\text{\times}{10}^{4}$ that satisfy (6). For each dataset, we compute the standard, unconditional MLE of $\beta$, $\hat{\beta}_{\text{MLE}}=(X^{\top}X)^{-1}X^{\top}Y$. For the same datasets, we compute the conditional MLE of $\beta_{1}$ that arises from numerically maximizing the conditional likelihood in (B.9). Each dataset corresponds to a point on the above scatterplots, with the standard point estimate on the horizontal axis and the conditional point estimate on the vertical axis. The point estimates in the left plot are computed using the true value of $\sigma^{2}$, and the point estimates in the right plot use the variance estimate $\tilde{\sigma}^{2}$} in (18).
B.2 Confidence intervals
We now consider constructing selective confidence intervals for $\beta_{j}$, i.e., confidence intervals that attain the nominal coverage, conditional on having rejected $H_{0}^{1:p}$. First, we extend our selective testing framework to accommodate more general hypotheses of the form $H_{0}^{j}(b):\beta_{j}=b$. With this in hand, confidence intervals can be constructed by means of test inversion, i.e., a level $\left(1-\alpha\right)$ confidence interval will contain all values of $b$ for which we would not reject $H_{0}^{j}(b)$ at level $\alpha$.
To test $H_{0}^{j}(b):\beta_{j}=b$, note that
${F_{H_{0}^{j}(b)}}:=(n-p-1)\cdot\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}$
unconditionally follows an $F_{1,n-p-1}$ distribution under $H_{0}^{j}(b)$ (Seber & Lee, 2003, §4.3). Thus the analogous selective p-value to that in (9) is
${p_{H_{0}^{j}(b)\mid E}}(y):=Pr_{\beta_{j}=b}\Bigg($ $\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)\Big|Y\in E(y)\Bigg),$ (B.10)
where
$r^{\prime}(y):=\frac{(y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(y-bX_{j})}{y^{\top}(I_{n-1}-P_{X})y}.$
Per Lemma E.1, the random vectors $(P_{X}-P_{{X}_{-j}})Y$, $(I_{n-1}-P_{X})Y$, and $P_{{X}_{-j}}Y$ are mutually independent. As in the proof of Proposition 2.1, this allows us to rewrite the selective p-value introduced in (B.10) as
${p_{H_{0}^{j}(b)\mid E}}(y)=Pr_{\beta_{j}=b}\Bigg($ $\frac{(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r^{\prime}(y)$
$\mid Y^{\top}(P_{X}-P_{X_{-j}})Y-cY^{\top}(I_{n-1}-P_{X})Y\geq d\Bigg),$ (B.11)
where $c$ and $d$ are defined in (11).
Next, we observe that
$(Y-bX_{j})^{\top}(P_{X}-P_{X_{-j}})(Y-bX_{j})$
$\hskip 14.22636pt=Y^{\top}(P_{X}-P_{X_{-j}})Y-2bX_{j}^{\top}(P_{X}-P_{X_{-j}})Y+b^{2}X_{j}^{\top}(P_{X}-P_{X_{-j}})X_{j}.$
Furthermore, from the proof of Lemma E.1, we have that
$(P_{X}-P_{X_{-j}})Y\sim\mathcal{N}_{n-1}(\beta_{j}(P_{X}-P_{X_{-j}})X_{j},\sigma^{2}(P_{X}-P_{X_{-j}})),$
and
$\frac{1}{\sigma^{2}}Y^{\top}(I_{n-1}-P_{X})Y\sim X^{2}_{n-p-1}.$
Therefore, $p_{H_{0}^{j}(b)\mid E}$} in (B.2) can be rewritten as
${p_{H_{0}^{j}(b)\mid E}}(y)=Pr\Bigg($ $\frac{R^{\top}R-2bX_{j}^{\top}R+b^{2}X_{j}^{\top}(P_{X}-P_{X_{-j}})X_{j}}{\sigma^{2}Z}\geq r^{\prime}(y)$
$\mid R^{\top}R-c\sigma^{2}Z\geq d\Bigg),$ (B.12)
where $Z\sim\chi^{2}_{n-p-1}$ and $R\sim\mathcal{N}_{n-1}(b(P_{X}-P_{X_{-j}})X_{j},\sigma^{2}(P_{X}-P_{X_{-j}}))$ with $Z$ independent of $R$. This can be readily approximated via Monte Carlo. Using this p-value, we define selective confidence intervals:
$\mathrm{CI}_{\alpha}^{j}(Y\mid Y\in E(y)):=\left\{b\in\mathbb{R}:{p_{H_{0}^{j}(b)\mid E}}\geq\alpha\right\}.$ (B.13)
The following proposition is corroborated in the left panels of Fig. 6 and Fig. C.2.
Proposition B.2.
For any $\alpha\in(0,1)$, $\mathrm{CI}_{\alpha}^{j}(Y\mid Y\in E(y))$ in (B.13) provides nominal coverage for the parameter $\beta_{j}$ conditional on having rejected $H_{0}^{1:p}$ in (2). That is,
$\Pr\left(\beta_{j}\in\mathrm{CI}_{\alpha}^{j}(Y\mid Y\in E(y))\mid Y\in E_{1}\right)=1-\alpha$ (B.14)
for $E_{1}$ defined in (6).
Proof of Proposition B.2.
First we note that under the null $H_{0}^{j}(b):\beta_{j}=b$, the p-value $p_{H_{0}^{j}(b)\mid E}$} in (B.2) follows a $\operatorname{Unif}(0,1)$ distribution, conditional on $Y\in E_{1}$ (6). This follows from a very similar argument to that in the proof of Theorem 2.1. Hence,
$\Pr\left(\beta_{j}\in\mathrm{CI}_{\alpha}^{j}(Y\mid Y\in E(y))\mid Y\in E_{1}\right)$ $=\Pr\left(p_{H_{0}^{j}(b)\mid E}>\alpha\mid Y\in E_{1}\right)$
$=1-\Pr\left(p_{H_{0}^{j}(b)\mid E}\leq\alpha\mid Y\in E_{1}\right)$
$=1-\alpha.$
Additionally, $p_{H_{0}^{j}(b)\mid E}$} relies on the value of $\sigma^{2}$. Therefore we define an analogous version of $p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}$} in (21) for the hypothesis $H_{0}^{j}(b):\beta_{j}=b$, i.e.,
${p^{\tilde{\sigma}^{2}}_{H_{0}^{j}(b)\mid E}}(y)=Pr\Bigg($ $\frac{R^{\top}R-2bX_{j}^{\top}R+b^{2}X_{j}^{\top}(P_{X}-P_{X_{-j}})X_{j}}{{\tilde{\sigma}^{2}}Z}\geq r^{\prime}(y)$
$\mid R^{\top}R-c{\tilde{\sigma}^{2}}Z\geq d\Bigg),$ (B.15)
where we now let $R\sim\mathcal{N}\left(b(P_{X}-P_{X_{-j}})X_{j},{\tilde{\sigma}^{2}}(P_{X}-P_{X_{-j}})\right)$. In practice, we use
$\mathrm{CI}_{\alpha}^{j,\tilde{\sigma}^{2}}(Y\mid Y\in E(y)):=\left\{b\in\mathbb{R}:{p^{\tilde{\sigma}^{2}}_{H_{0}^{j}(b)\mid E}}\geq\alpha\right\}$ (B.16)
as our confidence interval.
Remark 4.
The expression for $p_{H_{0}^{j}(b)\mid E}$} in (B.2) can be re-written for greater computational efficiency. Specifically, noting that $P_{X}-P_{X_{-j}}$ is a rank one projection matrix, let $U_{j}\in\mathbb{R}^{n}$ satisfy $U_{j}^{\top}U_{j}=1$ and $U_{j}U_{j}^{\top}=P_{X}-P_{X_{-j}}.$ Then
$Y^{\top}(P_{X}-P_{X_{-j}})Y=\|U_{j}^{\top}Y\|_{2}^{2}.$
Thus, noting that $U_{j}^{\top}Y\sim\mathcal{N}(U_{j}^{\top}X_{j}\beta_{j},\sigma^{2}),$ it follows from (B.2) that
${p_{H_{0}^{j}(b)\mid E}}=Pr_{\beta_{j}=b}\Bigg($ $\frac{\|U_{j}^{\top}Y\|_{2}^{2}-2b(X_{j}^{\top}U_{j})(U_{j}^{\top}Y)+b^{2}(X_{j}^{\top}U_{j})^{2}}{\sigma^{2}Z}\geq r^{\prime}(y)$
$\mid\|U_{j}^{\top}Y\|^{2}-c\sigma^{2}Z\geq d\Bigg)$
$=Pr\Bigg($ $\frac{G^{2}-2b(X_{j}^{\top}U_{j})G+b^{2}(X_{j}^{\top}U_{j})^{2}}{\sigma^{2}Z}\geq r^{\prime}(y)\mid G^{2}-c\sigma^{2}Z\geq d\Bigg),$ (B.17)
where $Z\sim\chi^{2}_{n-p-1}$ and $G\sim\mathcal{N}(U_{j}^{\top}X_{j}b,\sigma^{2}).$ The advantage of (4) over (B.2) is that the former can be computed using only univariate variables, whereas the latter involves the generation of (n1)(n-1)-dimensional random vectors in each Monte Carlo replicate.
We observe that the p-value in (B.2) depends on $X_{j}$ in a way that cannot be deduced from statistics such as $\mathrm{R}^{2}$, $\mathrm{RSE}$ (22), and $F_{H_{0}^{M}}$ (4). Thus confidence intervals cannot be computed retrospectively, in the sense of Section 3.
Appendix C Extended simulation results
In the setting of Section 5, we modify the simulation so that $\beta_{j}=0.1$ for $j>1$. The results are shown in Fig. C.1 and Fig. C.2.
Figure C.1: Details are as in Fig. 5, but with $\beta_{j}=0.1$ for $j>1$.
Compared to the setting in Section 5 and Fig. 5, in the left panel of Fig. C.1 the standard p-value $p_{H_{0}^{M}}$ in (5) is less anti-conservative, and in the left panel of Fig. C.2 the standard confidence intervals achieve coverage closer to the nominal level, conditional on rejection of $H_{0}^{1:p}$ (2). These improvements can be attributed to the fact that rejection of $H_{0}^{1:p}$ is highly likely when $\beta_{2}=\ldots=\beta_{p}=0.1$, so conditioning on $Y\in E$ (8) has relatively little impact on the test of $H_{0}^{1}$. Thus, the performance of standard inference methods approaches the performance of selective inference methods.
The middle panel of Fig. C.1 shows that when $\beta_{2}=\cdots=\beta_{p}=0.1$, the overall test of $H_{0}^{1:p}$ based on the full dataset has higher power than a test based only on the training data; both tests have higher power to reject $H_{0}^{1:p}$ than in the setting of Fig. 5, in which $\beta_{2}=\ldots=\beta_{p}=0$. Comparing the right panel of Fig. C.1 to Fig. 5 reveals that the selective test of $H_{0}^{1}$, which accounts for the rejection of $H_{0}^{1:p}$ and uses the full dataset, has greater power when $\beta_{2}=\cdots=\beta_{p}=0.1$ than when $\beta_{2}=\ldots=\beta_{p}=0$. This is because in the former case, the rejection of $H_{0}^{1:p}$ is largely driven by the signal in the coefficients $\beta_{j}$ for $j>1$, meaning that substantial signal about $\beta_{1}$ remains available for inference.
The middle and right-hand plot of Fig. C.2 exhibit similar patterns to those in Fig. 6, so we do not provide further commentary.
Figure C.2: Details are as in Fig. 6, but with $\beta_{j}=0.1$ for $j>1$.
Appendix D Geometric intuition
In this section we explain the geometric view of the F-screening procedure displayed in Fig. 2 and Fig. 3. For simplicity, we assume that $\beta_{0}=0$ in the context of (1). We let $\alpha_{0}=\alpha=0.05.$
D.1 Figure 3
Let $Y=\begin{bmatrix}Y_{1}\\ Y_{2}\end{bmatrix}$, $X=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, and $\sigma^{2}=1.$ In this setting, we assume that the error variance is known, and so we conduct chi-squared tests instead of $F$-tests.
  • Step 1 of Box 1: Under $H_{0}^{1:p}:\beta_{1}=\beta_{2}=0$, we have that
    $\frac{Y^{T}P_{X}Y}{\sigma^{2}}=Y^{T}Y=Y_{1}^{2}+Y_{2}^{2}\sim\chi^{2}_{2}.$
    Thus, if the observed data $(y_{1},y_{2})$ lies outside the circle defined by $Y_{1}^{2}+Y_{2}^{2}=F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})$, then we reject this overall test, where $F_{\chi^{2}_{p}}$ is the CDF of a $\chi^{2}_{p}$ random variable. This corresponds to the circle in both plots of Fig. 3.
  • Step 2 of Box 1 (standard inference, left plot of Figure 3): Under $H_{0}^{1}:\beta_{1}=0$, we have that
    $\frac{Y^{T}(P_{X}-P_{X_{-1}})Y}{\sigma^{2}}$ $=Y^{T}\left(\begin{bmatrix}1&0\\ 0&1\end{bmatrix}-\begin{bmatrix}0\\ 1\end{bmatrix}\left(\begin{bmatrix}0&1\end{bmatrix}\begin{bmatrix}0\\ 1\end{bmatrix}\right)^{-1}\begin{bmatrix}0&1\end{bmatrix}\right)Y$
    $=Y_{1}^{2}\sim\chi_{1}^{2},$
    unconditional on the rejection of $H_{0}^{1:p}.$ Thus, if the observed data $(y_{1},y_{2})$ satisfies $|y_{1}|\geq\sqrt{F_{\chi^{2}_{1}}^{-1}(1-\alpha)}$, then we reject this standard test of $H_{0}^{1}$. This corresponds to the vertical lines in the left plot of Fig. 3.
  • Step 2 of Box 1 (selective inference, right plot in Figure 3): We construct a selective p-value for $H_{0}^{1}$ based on chi-square tests in a manner similar to Proposition 2.1. We arrive at the p-value
    $p_{\chi}(y):=\Pr_{\beta_{1}=0}\Big(Y^{\top}(P_{X}-P_{X_{-1}})Y\geq\sigma^{2}y^{\top}(P_{X}-P_{X_{-1}})y$
    $\quad\quad\quad\quad\quad\quad\mid Y^{\top}(P_{X}-P_{X_{-1}})Y\geq\sigma^{2}F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})-y^{\top}P_{X_{-1}}y\Big)$
    $=\Pr\left(W\geq\sigma^{2}y^{\top}(P_{X}-P_{X_{-1}})y\mid W\geq\sigma^{2}F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})-y^{\top}P_{X_{-1}}y\right)$
    $=\min\left\{1,\frac{\Pr\left(W\geq\sigma^{2}y^{\top}(P_{X}-P_{X_{-1}})y\right)}{\Pr\left(W\geq\sigma^{2}F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})-y^{\top}P_{X_{-1}}y\right)}\right\}$
    $=\min\left\{1,\frac{1-F_{\chi_{1}^{2}}\left(\sigma^{2}y^{\top}(P_{X}-P_{X_{-1}})y\right)}{1-F_{\chi_{1}^{2}}\left(\sigma^{2}F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})-y^{\top}P_{X_{-1}}y\right)}\right\},$
    where $W\sim\chi^{2}_{1}$. In our setting, $\sigma^{2}y^{\top}(P_{X}-P_{X_{-1}})y=y_{1}^{2}$ and $y^{\top}P_{X_{-1}}y=y_{2}^{2}$. Thus, if the observed data $(y_{1},y_{2})$ lead to
    $p_{\chi}(y)=\min\left\{1,\frac{1-F_{\chi_{1}^{2}}\left(y_{1}^{2}\right)}{1-F_{\chi_{1}^{2}}\left(F_{\chi^{2}_{2}}^{-1}(1-\alpha_{0})-y_{1}^{2}\right)}\right\}\leq\alpha,$
    then we reject the selective test of $H_{0}^{1}.$ This corresponds to the purple region of the right plot in Fig. 3.
D.2 Figure 2
Now, we assume that $\sigma^{2}$ is unknown and
$Y=\begin{bmatrix}Y_{1}\\ Y_{2}\\ Y_{3}\end{bmatrix}\quad\text{and}\quad X=\begin{bmatrix}1&0\\ 0&1\\ 0&0\end{bmatrix}$.
Since $\sigma^{2}$ is unknown, we use $F$-statistics to conduct the hypothesis tests.
  • Step 1 of Box 1: Under $H_{0}^{1:p}:\beta_{1}=\beta_{2}=0$,
    $\frac{Y^{T}P_{X}Y}{Y^{T}(I_{n-1}-P_{X})Y}=\frac{Y_{1}^{2}+Y_{2}^{2}}{Y_{3}^{2}}\sim F_{2,1}.$
    Thus we reject $H_{0}^{1:p}$ if
    $\frac{y_{1}^{2}+y_{2}^{2}}{y_{3}^{2}}\geq F^{-1}_{2,1}(1-\alpha_{0}).$
    The rejection boundary corresponds to the double cone in both panels of Fig. 2.
  • Step 2 of Box 1: Under $H_{0}^{1}:\beta_{1}=0$,
    $\frac{Y^{T}(P_{X}-P_{X_{-1}})Y}{Y^{T}(I_{n-1}-P_{X})Y}$ $=\frac{Y_{1}^{2}}{Y_{3}^{2}}\sim F_{1,1},$
    unconditional on rejection of $H_{0}^{1:p}.$ Therefore, we reject $H_{0}^{1}$ if
    $\frac{y_{1}^{2}}{y_{3}^{2}}\geq F^{-1}_{1,1}(1-\alpha)=:b.$
    Thus the rejection boundary corresponds to the planes defined by $y_{1}/y_{3}\geq\sqrt{b}$ and $y_{1}/y_{3}\geq-\sqrt{b}$ in both panels of Fig. 2.
Appendix E Proofs of theoretical results in Sections 2 and 3
In this appendix, we provide proofs for all theorems and propositions from the main text. We begin by stating and proving a lemma that will be useful.
E.1 A useful lemma
Lemma E.1.
The following statements hold under the model $Y=X\beta+\epsilon$ where $X\in\mathbb{R}^{(n-1)\times p}$ and $\epsilon\sim\mathcal{N}_{n-1}\left(0,\sigma^{2}I_{n-1}\right)$:
  1. 1.
    Under $H_{0}^{M}:\beta_{M}=0$, we have that $\tfrac{1}{\sigma^{2}}Y^{\top}(P_{X}-P_{{X}_{-M}})Y\sim\chi_{m}^{2}$.
  2. 2.
    $\tfrac{1}{\sigma^{2}}Y^{\top}(I_{n-1}-P_{X})Y\sim\chi^{2}_{n-p-1}$.
  3. 3.
    In general, the distribution of $Y^{\top}P_{{X}_{-M}}Y$ involves $\beta_{j}$ for $j\notin M$.
  4. 4.
    $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$, $Y^{\top}(I_{n-1}-P_{X})Y$, and $Y^{\top}P_{{X}_{-M}}Y$ are mutually independent.
Proof of Lemma E.1.
We begin by proving the fourth statement of Lemma E.1. First, note that $P_{X}-P_{X_{-M}}$, $I_{n-1}-P_{X}$ and $P_{X_{-M}}$ are projection matrices, and thus they are symmetric and idempotent. As a result, it follows that
$Y^{\top}(P_{X}-P_{X_{-M}})Y=Y^{\top}(P_{X}-P_{X_{-M}})^{\top}(P_{X}-P_{X_{-M}})Y=\left\lVert(P_{X}-P_{X_{-M}})Y\right\rVert_{2}^{2}.$
Similarly,
$Y^{\top}(I_{n-1}-P_{X})Y=\left\lVert(I_{n-1}-P_{X})Y\right\rVert_{2}^{2}\hskip 8.53581pt\text{and}\hskip 8.53581ptY^{\top}P_{X_{-M}}Y=\left\lVert P_{X_{-M}}Y\right\rVert_{2}^{2}.$
Thus it suffices to show that $(P_{X}-P_{X_{-M}})Y,$ $(I_{n-1}-P_{X})Y,$ and $P_{X_{-M}}Y$ are mutually independent. Letting
$A=\begin{bmatrix}P_{X}-P_{X_{-M}}\\ I_{n-1}-P_{X}\\ P_{X_{-M}}\end{bmatrix}$,
observe that $AY\sim\mathcal{N}_{3(n-1)}\left(AX\beta,\sigma^{2}AA^{\top}\right)$. Since this is a (degenerate) multivariate normal distribution, proving that $(P_{X}-P_{X_{-M}})Y,$ $(I_{n-1}-P_{X})Y,$ and $P_{X_{-M}}Y$ are mutually independent is equivalent to proving that the off-diagonal blocks of the covariance matrix are all zero. The off-diagonal blocks of the covariance matrix, omitting the constant $\sigma^{2}$, are as follows:
$(P_{X}-P_{X_{-M}})(I_{n-1}-P_{X})=(P_{X}-P_{X_{-M}}-P_{X}+P_{X_{-M}}P_{X})=P_{X_{-M}}P_{X}-P_{X_{-M}}$
$(I_{n-1}-P_{X})P_{X_{-M}}=P_{X_{-M}}-P_{X}P_{X_{-M}}$
$(P_{X}-P_{X_{-M}})P_{X_{-M}}=P_{X}P_{X_{-M}}-P_{X_{-M}}.$
Furthermore, noting that $R(X_{-M})\subseteq R(X)$ where $R(\cdot)$ denotes the column space, it follows from properties of projection matrices that $P_{X}P_{X_{-M}}=P_{X_{-M}}$, and hence also that $P_{X_{-M}}P_{X}=P_{X_{-M}}$.
The fourth statement of Lemma E.1 follows.
To prove claims 1-3, we recall that if $Y\sim\mathcal{N}_{n-1}(\mu,\sigma^{2}I_{n-1})$ and $P$ a projection matrix of rank $r$, then $Y^{\top}PY/\sigma^{2}\sim\chi^{2}_{r}(\lambda)$ where $\lambda=\mu^{\top}P\mu/(2\sigma^{2})$ is the noncentrality parameter. Using this result, to prove part 1 of Lemma E.1, observe that $\frac{1}{\sigma}Y\sim\mathcal{N}_{n-1}\left(\frac{1}{\sigma}X\beta,I_{n-1}\right)$ and $I_{n-1}-P_{X}$ is a projection matrix of rank np1n-p-1. It then follows that
$\frac{1}{\sigma^{2}}Y^{\top}(I_{n-1}-P_{X})Y\sim\chi^{2}_{n-p-1}(0)\equiv\chi^{2}_{n-p-1}.$
To prove part 2, note that
$(P_{X}-P_{X_{-M}})X\beta$ $=(P_{X}-P_{X_{-M}})X_{M}\beta_{M}+(P_{X}-P_{X_{-M}})X_{-M}\beta_{-M}$
$=(P_{X}-P_{X_{-M}})X_{M}\beta_{M}$
$=P_{X}X_{M}\beta_{M}-P_{X_{-M}}X_{M}\beta_{M}$
$=X_{M}\beta_{M}-P_{X_{-M}}X_{M}\beta_{M}$
$=(I_{n-1}-P_{X_{-M}})X_{M}\beta_{M}.$
Since $(P_{X}-P_{X_{-M}})$ is a projection matrix of rank $\left\lvert M\right\rvert=m$, it follows that
$\frac{1}{\sigma^{2}}Y^{\top}(P_{X}-P_{X_{-M}})Y\sim\chi^{2}_{m}\left(\frac{(X_{M}\beta_{M})^{\top}(I_{n-1}-P_{X_{-M}})X_{M}\beta_{M}}{2\sigma^{2}}\right).$ (E.1)
Therefore, under the null $H_{0}^{M}:\beta_{M}=0$, $\frac{1}{\sigma^{2}}Y^{\top}(P_{X}-P_{X_{-M}})Y\sim\chi^{2}_{m}.$
Finally, to prove the third statement, note that $P_{X_{-M}}$ is a projection matrix of rank pmp-m. Hence,
$\frac{1}{\sigma^{2}}Y^{\top}P_{X_{-M}}Y\sim\chi^{2}_{p-m}\left(\frac{1}{2\sigma^{2}}(X\beta)^{\top}P_{X_{-M}}X\beta\right).$ (E.2)
Unless $X_{M}^{\top}X_{M}=0$, the non-centrality parameter involves both $\beta_{M}$ and $\beta_{-M}.$
E.2 Proof of Theorem 2.1
We begin with the following remark, which can be viewed as an extension of the probability integral transform to accommodate conditioning.
Remark E.1.
Let $T$ be a random variable and $C$ an event. Let
$F_{T\mid C}(t)=\Pr(T\leq t\mid C),$
i.e., the conditional CDF of $T$ given $C$. Define $D=F_{T\mid C}(T)$. Then
$\Pr(D\leq\alpha\mid C)$ $=\Pr(F_{T\mid C}(T)\leq\alpha\mid C)$
$=\Pr(T\leq F_{T\mid C}^{-1}(\alpha)\mid C)$
$=F_{T\mid C}(F_{T\mid C}^{-1}(\alpha))$
$=\alpha.$
Thus $D\sim\operatorname{Unif}\left(0,1\right)$.
Proof of Theorem 2.1.
Remark E.1 implies that, for $\alpha^{\prime}\in(0,1)$ and $E$ defined in (8),
$Pr_{\beta_{M}=0}({p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\mid Y\in E)=\alpha^{\prime}.$
Equivalently,
$\mathbb{E}\left[\mathbbm{1}\{{p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\}\mid Y\in E\right]=\alpha^{\prime}.$
Because $Y\in E\implies Y\in E_{1}$ for $E_{1}$ defined in (6) and $E(y)$ defined in (8), when we apply the law of total expectation, we obtain
$Pr_{\beta_{M}=0}$ $({p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\mid Y\in E_{1})$
$=\mathbb{E}\left[\bm{I}\{{p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\}\mid Y\in E_{1}\right]$
$=\mathbb{E}\Big[\mathbb{E}\left[\bm{I}\{{p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\}\mid Y\in E_{1}\cap E_{2}(y)\right]\mid Y\in E_{1}\Big]$
$=\mathbb{E}\Big[\mathbb{E}\left[\bm{I}\{{p_{H_{0}^{M}\mid E}}\leq\alpha^{\prime}\}\mid Y\in E\right]\mid Y\in E_{1}\Big]$
$=\mathbb{E}[\alpha^{\prime}\mid Y\in E_{1}]$
$=\alpha^{\prime}.$
E.3 Proof of Proposition 2.1
We begin with a lemma.
Lemma E.2.
For $p_{H_{0}^{M}\mid E}$} in (9), it holds that
${p_{H_{0}^{M}\mid E}}(y)=\Pr_{\beta_{M}=0}$ $\Bigg(\frac{Y^{\top}(P_{X}-P_{X_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)$
$\mid Y^{\top}\left(P_{X}-P_{X_{-M}}\right)Y-cY^{\top}\left(I_{n-1}-P_{X}\right)Y\geq d\Bigg),$ (E.3)
where $c$, $d$ and $r(y)$ are defined in (11).
Proof of Lemma E.2.
Note that, with $E_{1}$ defined in (6) and $c$ defined in (11),
$Y\in E_{1}$ $\iff\frac{n-p-1}{p}\frac{Y^{\top}P_{X}Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq F^{-1}_{p,n-p-1}(1-\alpha)$
$\iff Y^{\top}P_{X}Y-cY^{\top}(I_{n-1}-P_{X})Y\geq 0$
$\iff Y^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y\geq-Y^{\top}P_{X_{-M}}Y.$
From part 4 of Lemma E.1, we have that $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$, $Y^{\top}(I_{n-1}-P_{X})Y$, and $Y^{\top}P_{{X}_{-M}}Y$ are mutually independent. Importantly, this implies that $Y^{\top}P_{{X}_{-M}}Y$ is independent of any function of $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$. Therefore $Y^{\top}P_{{X}_{-M}}Y$ is independent of both $Y^{\top}(P_{X}-P_{X_{-M}})Y/Y^{\top}(I_{n-1}-P_{X})Y$ and $Y^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$, as well as their intersection. For $d$ defined as in (11), it follows that
${p_{H_{0}^{M}\mid E}}(y)$ $=\Pr_{\beta_{M}=0}\Bigg(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)\mid Y\in E(y)\Bigg)$
$=\Pr_{\beta_{M}=0}\Bigg(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)\mid Y\in E_{2}(y),\hskip 5.69054ptY\in E_{1}\Bigg)$
$=\Pr_{\beta_{M}=0}\Bigg(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)\mid Y^{\top}P_{{X}_{-M}}Y=-d,$
$\hskip 72.26999ptY^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y\geq-Y^{\top}P_{X_{-M}}Y\Bigg)$
$=Pr_{\beta_{M}=0}\Bigg(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)\mid Y^{\top}P_{{X}_{-M}}Y=-d,$
$\hskip 72.26999ptY^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y\geq d\Bigg)$
$=Pr_{\beta_{M}=0}\Bigg(\frac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq r(y)\mid$
$\hskip 72.26999ptY^{\top}(P_{X}-P_{X_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y\geq d\Bigg),$
where the final equality follows from the fact that if $A,B,$ and $C$ are random events where $(A,B)$ is independent of $C$ and $B$ is independent of $C$, then $\Pr(A|B,C)=\Pr(A|B).$
Now we can prove Proposition 2.1.
Proof of Proposition 2.1.
By Lemma E.1 parts 1 and 2, under $H_{0}^{1:p}$, $\tfrac{Y^{\top}(P_{X}-P_{{X}_{-M}})Y}{Y^{\top}(I_{n-1}-P_{X})Y}$ is distributed as $W/Z$ and $Y^{\top}(P_{X}-P_{{X}_{-M}})Y-cY^{\top}(I_{n-1}-P_{X})Y$ is distributed as $\sigma^{2}W-c\sigma^{2}Z$ where $W\sim\chi_{m}^{2}$ and $Z\sim\chi_{n-p-1}^{2}$. Furthermore, by part 4 of Lemma E.1, $Y^{\top}(P_{X}-P_{{X}_{-M}})Y$ and $Y^{\top}(I_{n-1}-P_{X})Y$ are independent. Using the representation of $p_{H_{0}^{M}\mid E}$} introduced in (E.3), the result follows. ∎
E.4 Proof of Proposition 2.2
We begin with two lemmas.
Lemma E.3.
Consider a fixed sequence of design matrices $X_{n}\in\mathbb{R}^{n\times p}$ and corresponding realizations of the linear model $Y_{n}=X_{n}\beta+\epsilon_{n}$ from (1). Suppose that $\frac{1}{n}X_{n}^{\top}X_{n}\to Q$ for some positive definite $Q$. Then $\exists\delta>0$ and $N>0$ such that $\Pr(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2})\geq\delta$ for all $n>N$, where $W\sim\chi^{2}_{m}$, $Z_{n}\sim\chi^{2}_{n-p-1}$, $c_{n}=\frac{p}{n-p-1}F^{-1}_{p,n-p-1}\left(1-\alpha_{0}\right)$, and $d(Y_{n})=-Y_{n}^{\top}P_{\left(X_{n}\right)_{-M}}Y_{n}$, with $W$, $Z_{n}$, and $d(Y_{n})$ independent.
Proof of Lemma E.3.
In what follows, we suppress subscripts on $X$ involving $n$ to enhance readability.
To show that $\Pr(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2})>0$ as $n\to\infty$, first observe that $c_{n}\cdot Z_{n}$ converges in probability to $(\chi^{2}_{p})^{-1}(1-\alpha_{0})$ as $n\to\infty$. On the other hand, $d(Y_{n})$ is a scaled chi-square random variable with pmp-m degrees of freedom and non-centrality parameter $\tfrac{1}{2\sigma^{2}}(X\beta)^{\top}P_{X_{-M}}(X\beta)$. We treat the cases where $\beta\neq 0$ and $\beta=0$ separately.
Case 1: Suppose that $\beta\neq 0$. Then the non-centrality parameter of $d(Y_{n})$ is possibly non-zero. By our assumption that $\frac{1}{n}X^{\top}X\to Q$ deterministically for some positive definite matrix $Q\in\mathbb{R}^{p\times p}$, it follows that $\frac{1}{n}X^{\top}X_{-M}\to Q_{\cdot,-M}$ and $\frac{1}{n}X_{-M}^{\top}X_{-M}\to Q_{-M,-M}$ for $Q_{\cdot,-M}\in\mathbb{R}^{p\times(p-m)}$ and positive definite $Q_{-M,-M}\in\mathbb{R}^{(p-m)\times(p-m)}$. Hence,
$\tfrac{1}{n}\beta^{\top}X^{\top}P_{X_{-M}}X\beta$ $=\beta^{\top}\left(\tfrac{1}{n}X^{\top}X_{-M}\right)\left(\tfrac{1}{n}X_{-M}^{\top}X_{-M}\right)^{-1}\left(\tfrac{1}{n}X_{-M}^{\top}X\right)\beta$
$\to\beta^{\top}Q_{\cdot,-M}(Q_{-M,-M})^{-1}Q_{\cdot,-M}^{\top}\beta.$
Thus, when $\beta\neq 0$, $d(Y_{n})$ is distributed as a scaled chi-square random variable with non-centrality parameter that is $O(n)$. The sequence of expectations of $d(Y_{n})/n$ converges and its variance is $O(1/n).$ Thus by Chebyshev’s inequality, it follows that $\phi_{n}:=d(Y_{n})/n\overset{P}{\to}\phi^{*}$ for some fixed, negative $\phi^{*}$, where $\overset{P}{\to}$ denotes convergence in probability.
Therefore, for fixed $\varepsilon>0$ and $\eta>0$, we have that $\Pr(|\phi_{n}-\phi^{*}|\geq\varepsilon)<\eta$ and $\Pr(|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha)|\geq\varepsilon)<\eta$ for sufficiently large $n$. For such an $n$, define
$G_{n}(\varepsilon)=\{(\phi_{n},Z_{n}):|\phi_{n}-\phi^{*}|<\varepsilon,\;|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha)|<\varepsilon\},$
so that $\Pr(G_{n}^{c}(\varepsilon))\leq 2\eta$. Hence
$\Pr\bigl(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\bigr)$ $=\Pr\bigl(W-c_{n}Z_{n}\geq n\phi_{n}/\sigma^{2}\bigr)$
$\geq\Pr\bigl(W\geq c_{n}Z_{n}+n\phi_{n}/\sigma^{2},\;G_{n}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2},\;G_{n}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2}\bigr)-\Pr\bigl(G_{n}^{c}(\varepsilon)\bigr)$
$\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi^{*}+\varepsilon)/\sigma^{2}\bigr)-2\eta.$
Note that $W\sim\chi^{2}_{m}$ has support $0,\infty)$ and, recalling that $\phi^{*}$ is negative, $(\chi^{2}_{p})^{-1}(1-\alpha)+\varepsilon+n(\phi_{n}^{*}+\varepsilon)/\sigma^{2}$ approaches -$\infty$ as $n\to\infty$. Since $\eta$ can be arbitrarily close to 0, we conclude that $\Pr\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)\to 1$ as $n\to\infty$, and hence is bounded away from 0 in the limit.
Case 2: Under the global null $H_{0}^{1:p}:\beta=0,$ the non-centrality parameter of $d(Y_{n})$ is 0. In this case, the distribution of $d(Y_{n})$ does not change with $n$. For fixed $\eta>0$, $\varepsilon>0$, we have that $\Pr(|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha_{0})|\geq\varepsilon)<\eta$ for sufficiently large $n$, and we can choose $B_{\eta}$ such that $\Pr(|d(Y_{n})|\geq B_{\eta})<\eta$. Then define
$G_{n}(\varepsilon,B_{\eta})=\{(d(Y_{n}),Z_{n}):|d(Y_{n})|<B_{\eta},\;|c_{n}Z_{n}-(\chi^{2}_{p})^{-1}(1-\alpha_{0})|<\varepsilon\},$
so that $\Pr(G_{n}^{c}(\varepsilon,B_{\eta}))\leq 2\eta$. Hence, similar to case 1,
$\Pr\bigl(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\bigr)\geq\Pr\bigl(W\geq(\chi^{2}_{p})^{-1}(1-\alpha_{0})+\varepsilon+B_{\eta}/\sigma^{2}\bigr)-2\eta.$
Letting $\eta$ approach 0, we conclude that $\Pr\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)>0$ in the limit since $(\chi^{2}_{p})^{-1}(1-\alpha_{0})+\varepsilon+B_{\eta}/\sigma^{2}$ is a fixed finite quantity that does not change (and in particular does not grow) with $n$.
Lemma E.4.
Suppose $f:\mathbb{R}^{n}\to\mathbb{R}$ is a function that satisfies $\Pr(f(Y_{n})>\varepsilon)\to 0$ for every $\varepsilon>0$. Then $\Pr(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1})\to 0$ for every $\varepsilon>0$.
Proof of Lemma E.4.
Note that $\Pr_{Y_{n}}(Y_{n}\in E_{1})\geq\alpha_{0}$ where $\alpha_{0}>0$ is the level of the F-test of $H_{0}^{1:p}$ in Step 1 of Box 1. Thus
$\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\in E_{1}\right)$
$\quad\quad\quad\quad\quad+\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\notin E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\notin E_{1}\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\Pr_{Y_{n}}\left(Y_{n}\in E_{1}\right)\to 0$
$\quad\implies\Pr_{Y_{n}}\left(f(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)\to 0.$
We now continue with the proof of Proposition 2.2.
Proof of Proposition 2.2.
In what follows, we use the notation $\Pr_{A}(\cdot)$ to denote probability taken with respect to a random variable $A$ in order to make explicit which quantity is treated as random in each probability calculation.
A test is consistent if the probability of rejection under any alternative converges to 1. Thus the test of $H_{0}^{M}$ based on the selective p-value in (10) is consistent if and only if the selective p-value converges to 0 under any alternative. That is, we want to show that if $\beta\neq 0$, then $\forall\varepsilon>0$,
$\Pr_{Y_{n}}$ $\left({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon\mid Y_{n}\in E_{1}\right)$
$=\Pr_{Y_{n}}\left(\Pr_{W,Z_{n}}\left(\tfrac{W}{Z_{n}}>r(Y_{n})|W-c_{n}Z_{n}\geq\tfrac{d(Y_{n})}{\sigma^{2}}\right)>\varepsilon\mid Y_{n}\in E_{1}\right)\to 0.$
By Lemma E.4, if $\Pr_{Y_{n}}({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon)\to 0$ then $\Pr_{Y_{n}}({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon\mid Y_{n}\in E_{1})\to 0$. Thus it is sufficient to show that whenever $\beta\neq 0,$
$\Pr_{Y_{n}}\left({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon\right)=\Pr\left(\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})|W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)>\varepsilon\right)\to 0.$
We begin by showing that $\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)\overset{P}{\to}0$ when $\beta_{M}\neq 0$, i.e. that
$\lim_{n\to\infty}\Pr_{Y_{n}}\left(\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)\right)=0.$
We have assumed that $\frac{1}{n}X^{\top}X\to Q$ deterministically for some positive definite matrix $Q\in\mathbb{R}^{p\times p}$, and it therefore follows that $\frac{1}{n}X_{M}^{\top}X_{-M}\to Q_{M,-M}$, $\frac{1}{n}X_{-M}^{\top}X_{-M}\to Q_{-M,-M}$, and $\frac{1}{n}X_{M}^{\top}X_{M}\to Q_{M,M}$ for $Q_{M,-M}\in\mathbb{R}^{m\times(p-m)}$ and positive definite matrices $Q_{M,M}\in\mathbb{R}^{m\times m}$ and $Q_{-M,-M}\in\mathbb{R}^{(p-m)\times(p-m)}$. Thus
$\tfrac{1}{n}X_{M}^{\top}(I-P_{X_{-M}})X_{M}$ $=\tfrac{1}{n}X_{M}^{\top}X_{M}-\tfrac{1}{n}X_{M}^{\top}P_{X_{-M}}X_{M}$
$=\tfrac{1}{n}X_{M}^{\top}X_{M}-\tfrac{1}{n}X_{M}^{\top}X_{-M}(X_{-M}^{\top}X_{-M})^{-1}X_{-M}^{\top}X_{M}$
$=\tfrac{1}{n}X_{M}^{\top}X_{M}-\left(\tfrac{1}{n}X_{M}^{\top}X_{-M}\right)\left(\tfrac{1}{n}X_{-M}^{\top}X_{-M}\right)^{-1}\left(\tfrac{1}{n}X_{-M}^{\top}X_{M}\right)$
$\to Q_{M,M}-Q_{M,-M}(Q_{-M,-M})^{-1}Q_{M,-M}^{\top}.$
It follows that $\frac{1}{2\sigma^{2}}(X_{M}\beta_{M})^{\top}(I-P_{X_{-M}})X_{M}\beta_{M}=O(n)$ when $\beta_{M}\neq 0$. Hence, by (E.1), when $\beta\neq 0$, $\frac{1}{\sigma^{2}}Y_{n}^{\top}(P_{X}-P_{X_{-M}})Y_{n}$ is distributed as a chi-square random variable with a non-centrality parameter that is $O(n).$ The sequence of expectations of $\frac{1}{n}Y_{n}^{\top}(P_{X}-P_{X_{-M}})Y_{n}$ converges, and the variance is $O(1/n)$. Hence, by Chebyshev’s inequality, it follows that $\frac{1}{n}Y_{n}^{\top}(P_{X}-P_{X_{-M}})Y_{n}\overset{P}{\to}C$ for some constant $C>0$. Additionally observe that $\frac{1}{n}Y_{n}^{\top}(I-P_{X})Y_{n}\overset{P}{\to}\sigma^{2}$, again by Chebyshev’s inequality. Hence, by the continuous mapping theorem
$r(Y_{n})=\frac{Y_{n}^{\top}(P_{X}-P_{{X}_{-M}})Y_{n}}{Y_{n}^{\top}(I_{n-1}-P_{X})Y_{n}}=\frac{\tfrac{1}{n}Y_{n}^{\top}(P_{X}-P_{{X}_{-M}})Y_{n}}{\tfrac{1}{n}Y_{n}^{\top}(I_{n-1}-P_{X})Y_{n}}\overset{P}{\to}\frac{C}{\sigma^{2}}>0.$
Further, $W/Z_{n}=\left(\tfrac{1}{n}W\right)/\left(\tfrac{1}{n}Z_{n}\right)\overset{P}{\to}0/1=0.$ That is, for any $\varepsilon>0,$ $\Pr_{W,Z_{n}}(W/Z_{n}>\varepsilon)\to 0.$ We conclude that $\Pr\left(W/Z_{n}>r(Y_{n})\right)\overset{P}{\to}0$.
Next, note that for any $y\in\mathbb{R}^{n}$,
$\Pr_{W,Z_{n}}\left(\frac{W}{Z_{n}}>r(y)|W-c_{n}Z_{n}\geq d(y)/\sigma^{2}\right)$ $=\frac{\Pr_{W,Z_{n}}\left(W/Z_{n}>r(y),\;W-c_{n}Z_{n}\geq d(y)/\sigma^{2}\right)}{\Pr_{W,Z_{n}}\left(W-c_{n}Z_{n}\geq d(y)/\sigma^{2}\right)}$
$\leq\frac{\Pr_{W,Z_{n}}\left(W/Z_{n}>r(y)\right)}{\Pr_{W,Z_{n}}\left(W-c_{n}Z_{n}\geq d(y)/\sigma^{2}\right)}.$
Thus,
$\Pr_{Y_{n}}$ $\left(\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})|W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)>\varepsilon\right)$
$\hskip 85.35826pt\leq\Pr_{Y_{n}}\left(\frac{\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)}{\Pr_{W,Z_{n}}\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)}>\varepsilon\right).$
Since $\Pr\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)$ is bounded away from zero in the limit by Lemma E.3, we conclude that
$\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)\overset{P}{\to}0$ $\implies\frac{\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)}{\Pr_{W,Z_{n}}\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)}\overset{P}{\to}0$
$\implies\Pr_{Y_{n}}\left(\frac{\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})\right)}{\Pr_{W,Z_{n}}\left(W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)}>\varepsilon\right)\to 0$
$\implies\Pr_{Y_{n}}\left(\Pr_{W,Z_{n}}\left(W/Z_{n}>r(Y_{n})|W-c_{n}Z_{n}\geq d(Y_{n})/\sigma^{2}\right)>\varepsilon\right)\to 0.$
Thus we have established that $\Pr_{Y_{n}}\left({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon\right)\to 0.$ By Lemma E.4, this implies that
$\Pr\left({p_{H_{0}^{M}\mid E}}(Y_{n})>\varepsilon\mid Y\in E_{1}\right)\to 0,$
and equivalently that $\Pr\left({p_{H_{0}^{M}\mid E}}(Y_{n})<\varepsilon\mid Y\in E_{1}\right)\to 1.$ This completes the proof.
E.5 Proof of Proposition 2.3
Proof of Proposition 2.3.
Recall that for a proportion $\rho\in[0,1]$ and a realization $\tilde{y}\in\mathbb{R}^{n}$ we define
$R_{\rho}(X)$ $=\{Y\in\mathbb{R}^{(\rho n)}:Y^{\top}P_{X}Y\geq c_{\chi}\},$
$R(\tilde{y};X)$ $=\{Y\in\mathbb{R}^{n}:Y^{\top}P_{X}Y\geq c_{\chi},\;P_{X_{-1}}Y=P_{X_{-1}}\tilde{y}\}$,
$c_{\chi}$ $=\sigma^{2}(\chi_{p}^{2})^{-1}(1-\alpha_{0}),$
$e_{\tilde{y}}$ $=c_{\chi}-\tilde{y}^{\top}P_{X_{-1}}\tilde{y}$.
Let $(X^{\mathrm{tr}},Y^{\mathrm{tr}})$ be the training dataset with $\rho$ proportion of the observations and $(X^{\mathrm{te}},Y^{\mathrm{te}})$ the test dataset with $(1-\rho)$ proportion of the observations. We begin by deriving the leftover Fisher information in $Y^{\mathrm{te}}$ about $\beta_{1}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$, where we treat $\beta_{2},\ldots,\beta_{p}$ as nuisance parameters. That is, we will find an expression for
$\left(\left[\left[\mathcal{I}_{Y^{\mathrm{te}}\mid Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})}(\beta;Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}}))\right]^{-1}\right]_{1,1}\right)^{-1},$ (E.4)
where $\mathcal{I}_{Y^{\mathrm{te}}\mid Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})}(\beta;Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}}))$ is the Fisher information matrix for the entire parameter vector $\beta$. We consider the inverse of the $(1,1)$ element of the inverted Fisher information matrix in order to isolate the information in $Y^{\mathrm{te}}$ about $\beta_{1}$ while accounting for the fact that $\beta_{-1}$ is also unknown.
Since $Y^{\mathrm{tr}}$ and $Y^{\mathrm{te}}$ are independent, the density of $Y^{\mathrm{te}}$ conditional on $Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})$ is equivalent to the unconditional density of $Y^{\mathrm{te}}$, i.e.
$f_{Y^{\mathrm{te}}\mid Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})}(y)=f_{Y^{\mathrm{te}}}(y)=(2\pi\sigma^{2})^{-n/2}\exp\left\{\frac{1}{2\sigma^{2}}(y-X^{\mathrm{te}}\beta)^{\top}(y-X^{\mathrm{te}}\beta)\right\}.$
Since we are concerned with the derivative of this expression with respect to $\beta,$ we will now consider this function as a likelihood of $\beta$ with $y\in\mathbb{R}^{n}$ fixed. Then the Hessian of the log likelihood is
$\frac{d^{2}}{d(\beta_{1},\;\beta_{-1})d(\beta_{1},\;\beta_{-1})^{\top}}\log f_{Y^{\mathrm{te}}}(\beta;y)=-\frac{1}{\sigma^{2}}\begin{bmatrix}(X^{\mathrm{te}})_{1}^{\top}(X^{\mathrm{te}})_{1}&(X^{\mathrm{te}})_{1}^{\top}(X^{\mathrm{te}})_{-1}\\ (X^{\mathrm{te}})_{-1}^{\top}(X^{\mathrm{te}})_{1}&(X^{\mathrm{te}})_{-1}^{\top}(X^{\mathrm{te}})_{-1}\\ \end{bmatrix}$.
Note that this Hessian is non-random because the test design matrix $X^{\mathrm{te}}$ is fixed. Treating $\beta_{-1}$ as a nuisance parameter, we take the Schur complement to obtain the $(1,1)$ entry of the inverse Fisher information matrix and conclude that the leftover Fisher information is
$\left(\left[\left[\mathcal{I}_{Y^{\mathrm{te}}\mid Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}})}(\beta;Y^{\mathrm{tr}}\in R_{\rho}(X^{\mathrm{tr}}))\right]^{-1}\right]_{1,1}\right)^{-1}$
$\quad=\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(X^{\mathrm{te}})_{1}-[(X^{\mathrm{te}})_{1}^{\top}(X^{\mathrm{te}})_{-1}][(X^{\mathrm{te}})_{-1}^{\top}(X^{\mathrm{te}})_{-1}]^{-1}[(X^{\mathrm{te}})_{-1}^{\top}(X^{\mathrm{te}})_{1}]\right)$
$\quad=\frac{1}{\sigma^{2}}\left((X^{\mathrm{te}})_{1}^{\top}(I-P_{(X^{\mathrm{te}})_{-1}})(X^{\mathrm{te}})_{1}\right).$
Next, we derive an expression for $\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))$ for some realization $\tilde{y}\in\mathbb{R}^{n}$. Let
  1. 1.
    $U_{1}\in\mathbb{R}^{n\times(n-p)}$ such that $U_{1}U_{1}^{\top}=I_{n}-P_{X}$ and $U_{1}^{\top}U_{1}=I_{n-p}$,
  2. 2.
    $U_{2}\in\mathbb{R}^{n}$ such that $U_{2}U_{2}^{\top}=P_{X}-P_{X_{-1}}$ and $U_{2}^{\top}U_{2}=1$,
  3. 3.
    $U_{3}\in\mathbb{R}^{n\times(p-1)}$ such that $U_{3}U_{3}^{\top}=P_{X_{-1}}$ and $U_{3}^{\top}U_{3}=I_{p-1}$.
Furthermore, define
$\tilde{Z}=U_{1}^{\top}Y\in\mathbb{R}^{n-p},\quad\tilde{W}=U_{2}^{\top}Y\in\mathbb{R},\quad\tilde{V}=U_{3}^{\top}Y\in\mathbb{R}^{p-1},\quad U=\begin{bmatrix}U_{1}&U_{2}&U_{3}\end{bmatrix}$.
Then note that
$U^{\top}U$ $=\begin{bmatrix}U_{1}^{\top}\\ U_{2}^{\top}\\ U_{3}^{\top}\end{bmatrix}\begin{bmatrix}U_{1}&U_{2}&U_{3}\end{bmatrix}=\begin{bmatrix}U_{1}^{\top}U_{1}&0&0\\ 0&U_{2}^{\top}U_{2}&0\\ 0&0&U_{3}^{\top}U_{3}\end{bmatrix}=I_{n},$
$UU^{\top}$ $=\begin{bmatrix}U_{1}&U_{2}&U_{3}\end{bmatrix}\begin{bmatrix}U_{1}^{\top}\\ U_{2}^{\top}\\ U_{3}^{\top}\end{bmatrix}=U_{1}U_{1}^{\top}+U_{2}U_{2}^{\top}+U_{3}U_{3}^{\top}=I_{n},$
$Y$ $=UU^{\top}Y=U_{1}U_{1}^{\top}Y+U_{2}U_{2}^{\top}Y+U_{3}U_{3}^{\top}Y=U_{1}\tilde{Z}+U_{2}\tilde{W}+U_{3}\tilde{V}$,
$Y$ $\in R(\tilde{y};X)\iff\tilde{W}^{2}\geq e_{\tilde{y}}\text{ and }\tilde{V}=U_{3}^{\top}\tilde{y}$.
Similarly to (B.7), this allows us to write the conditional density of $Y$ given $Y\in R(\tilde{y};X)$ as
$f_{Y\mid Y\in R(\tilde{y};X)}(y)=\frac{f_{\tilde{Z}}(U_{1}^{\top}y)f_{\tilde{W}}(U_{2}^{\top}y)\bm{I}_{\{y\;:\;U_{3}^{\top}y=U_{3}^{\top}\tilde{y}\}}(y)\bm{I}_{\{y\;:\;(U_{2}^{\top}y)^{2}\geq e_{\tilde{y}}\}}(y)}{\Pr\left(\tilde{W}^{2}\geq e_{\tilde{y}}\right)}$.
We now switch to viewing this expression as a likelihood of $\beta$ for a fixed realization $y\in\mathbb{R}^{n},$ and thus the indicator $\bm{I}_{\{y\;:\;U_{3}^{\top}y=U_{3}^{\top}\tilde{y}\}}(y)$ must be satisfied. Note that the density $f_{\tilde{Z}}$ does not involve $\beta$. Then,
$f_{Y\mid Y\in R(\tilde{y};X)}(\beta;y)\propto\frac{f_{\tilde{W}}(\beta;U_{2}^{\top}y)\bm{I}_{\{y\;:\;(U_{2}^{\top}y)^{2}\geq e_{\tilde{y}}\}}(y)}{\Pr\left(\tilde{W}^{2}\geq e_{\tilde{y}}\right)}=f_{\tilde{W}\mid\tilde{W}^{2}\geq e_{\tilde{y}}}(\beta;U_{2}^{\top}y).$
Note that $\tilde{W}=U_{2}^{\top}Y\sim N(U_{2}^{\top}X\beta,\sigma^{2})$, or equivalently $\tilde{W}\sim N_{1}(U_{2}^{\top}X_{1}\beta_{1},\sigma^{2})$. Thus, letting $\tilde{w}=U_{2}^{\top}y$,
$f_{\tilde{W}|\tilde{W}^{2}\geq e_{\tilde{y}}}(\beta;\tilde{w})=\frac{\frac{1}{\sigma}\phi\left(\frac{\tilde{w}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)\bm{I}_{\{\tilde{w}^{2}\geq e_{\tilde{y}}\}}(\tilde{w})}{1+\bm{I}_{\{e_{\tilde{y}}>0\}}\left(-\Phi\left(\frac{\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)+\Phi\left(\frac{-\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)\right)},$
where $\phi$ is the density and $\Phi$ the CDF of a standard normal random variable and $\bm{I}_{\{e_{\tilde{y}}>0\}}$ is the indicator that $\tilde{y}$ is such that $e_{\tilde{y}}>0$. Noting that the above display is free of $\beta_{-1}$, we rewrite this in terms of $Y$ as
$f_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};y)\propto\frac{\frac{1}{\sigma}\phi\left(\frac{U_{2}^{\top}y-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)\bm{I}_{\{(U_{2}^{\top}y)^{2}\geq e_{\tilde{y}}\}}(y)}{1+\bm{I}_{\{e_{\tilde{y}}>0\}}\left(-\Phi\left(\frac{\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)+\Phi\left(\frac{-\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}\right)\right)}.$
Accordingly, we simply consider the second derivative of the log likelihood with respect to $\beta_{1}$ instead of the entire Hessian (with respect to $\beta)$.
After some calculus, we find that if
$\bm{I}_{\{y\;:\;U_{3}^{\top}y=U_{3}^{\top}\tilde{y}\}}(y)=1\;\text{ and }\;\bm{I}_{\{y\;:\;(U_{2}^{\top}y)^{2}\geq e_{\tilde{y}}\}}(y)=1,$ (E.5)
then the second derivative has the expression
$\frac{d^{2}}{(d\beta_{1})^{2}}$ $\log f_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};y)$
$=\frac{X_{1}^{\top}(P_{X}-P_{X_{-1}})X_{1}}{\sigma^{2}}\left[-1+\bm{I}_{\{e_{\tilde{y}}>0\}}\left(\frac{[\phi(a)-\phi(b)]^{2}}{[1-\Phi(b)+\Phi(a)]^{2}}+\frac{\phi(a)\cdot a-\phi(b)\cdot b}{1-\Phi(b)+\Phi(a)}\right)\right]$
where
$a=\frac{-\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma},\quad b=\frac{\sqrt{e_{\tilde{y}}}-U_{2}^{\top}X_{1}\beta_{1}}{\sigma}.$
This second derivative is non-random. Additionally, the indicators in (E.5) are equivalent to $Y\in\mathbb{R}(\tilde{y};X)$. Thus the leftover Fisher information about $\beta_{1}$ in $Y$ conditional on $Y\in\mathbb{R}(\tilde{y};X)$ is
$\mathcal{I}_{Y\mid Y\in R(\tilde{y};X)}(\beta_{1};Y\in R(\tilde{y};X))=\tfrac{X_{1}^{\top}(I-P_{X_{-1}})X_{1}}{\sigma^{2}}\left[1-\bm{I}_{\{e_{\tilde{y}}>0\}}\left(\tfrac{[\phi(a)-\phi(b)]^{2}}{[1-\Phi(b)+\Phi(a)]^{2}}+\tfrac{\phi(a)\cdot a-\phi(b)\cdot b}{1-\Phi(b)+\Phi(a)}\right)\right].$
E.6 Proof of Proposition 2.4
Proof of Proposition 2.4.
We will show that ${\tilde{\sigma}^{2}}(\lambda)$ in (20) with $\lambda=\frac{1}{2\sigma^{2}}(X\beta)^{\top}X\beta$ is unbiased for $\sigma^{2}$ conditional on $Y\in E_{1}$, where $E_{1}$ defined in (6). From Lemma E.1, recall that $Y^{\top}(I_{n-1}-P_{X})Y/\sigma^{2}\sim\chi^{2}_{n-p-1}$. Additionally, $Y^{\top}P_{X}Y/\sigma^{2}\sim\chi^{2}_{p}(\lambda)$. Let $Z\sim\chi^{2}_{n-p-1}$ and $A\sim\chi^{2}_{p}(\lambda)$. Then,
${\mathbb{E}}\left(Y^{\top}(I_{n-1}-P_{X})Y\mid Y\in E_{1}\right)$ $={\mathbb{E}}\left(Y^{\top}(I_{n-1}-P_{X})Y\mid\frac{Y^{\top}P_{X}Y}{Y^{\top}(I_{n-1}-P_{X})Y}\geq c\right)$
$={\mathbb{E}}\left(\sigma^{2}Z\mid\frac{A}{Z}\geq c\right)$
$=\sigma^{2}{\mathbb{E}}\left(Z\mid\frac{A}{Z}\geq c\right),$
for $c$ defined in (11). The result follows. Under $H_{0}^{1:p}$, $\lambda=0,$ and therefore ${\tilde{\sigma}^{2}}={\tilde{\sigma}^{2}}(0)$ is unbiased for $\sigma^{2}.$
E.7 Proofs of Proposition 2.5 and Proposition 2.6
Proof of Proposition 2.5.
Define $\gamma_{n}={\mathbb{E}}(Z|A\geq cZ)$ for $A\sim\chi^{2}_{p}$, $Z\sim\chi^{2}_{n-p-1}$, and $c$ from (11). Then
${\tilde{\sigma}^{2}}-\sigma^{2}\overset{d}{=}\sigma^{2}\left(\frac{Z}{\gamma_{n}}-1\right)=\sigma^{2}\left(\frac{n-p-1}{\gamma_{n}}-1\right)+\sigma^{2}\left(\frac{Z}{n-p-1}-1\right)\frac{n-p-1}{\gamma_{n}},$ (E.6)
where $\overset{d}{=}$ denotes distributional equivalence.
We will first show that $\frac{n-p-1}{\gamma_{n}}$ converges to 1. Note that
$\gamma_{n}={\mathbb{E}}[Z|A\geq cZ]$ $=\frac{{\mathbb{E}}[Z\mathbbm{1}\{A\geq cZ\}]}{\Pr(A\geq cZ)}$
$=\frac{\mathrm{Cov}(Z,\mathbbm{1}\{A\geq cZ\})}{\Pr(A\geq cZ)}+\frac{{\mathbb{E}}[Z]{\mathbb{E}}[\mathbbm{1}\{A\geq cZ\}]}{\Pr(A\geq cZ)}$
$=\frac{\mathrm{Cov}(Z,\mathbbm{1}\{A\geq cZ\})}{\Pr(A\geq cZ)}+E(Z).$
Additionally, recalling that $c\cdot Z$ converges to a constant $(\chi^{2}_{p})^{-1}(1-\alpha_{0})$, $\Pr(A\geq cZ)$ is bounded away from 0 as $n\to\infty$. Hence,
$\frac{|\mathrm{Cov}(Z,\mathbbm{1}\{A\geq cZ\})|}{\Pr(A\geq cZ)}$ $\leq\frac{\sqrt{\mathrm{Var}(Z)\mathrm{Var}(\mathbbm{1}\{A\geq cZ\})}}{\Pr(A\geq cZ)}$
$=\frac{\sqrt{2(n-p-1)\Pr(A\geq cZ)(1-\Pr(A\geq c_{n}Z))}}{\Pr(A\geq cZ)}$
$\leq\frac{\sqrt{2(n-p-1)\cdot 1/4}}{\Pr(A\geq cZ)}$
$=O(\sqrt{n-p-1}),$
where the first line follows from Cauchy-Schwarz, and the second and third lines follow from the fact that the indicator is a Bernoulli random variable. Therefore, recalling that E(Z)=np1E(Z)=n-p-1, it follows that
$\gamma_{n}={\mathbb{E}}[Z|A\geq cZ]=n-p-1+O(\sqrt{n-p-1})=(n-p-1)(1+o(1)).$
Thus $\frac{n-p-1}{\gamma_{n}}$ converges to 1, from which we conclude the first term in (E.6) satisfies
$\sigma^{2}\left(\frac{n-p-1}{\gamma_{n}}-1\right)\overset{P}{\to}0.$
Turning to the second term in (E.6), note that $\left(\frac{Z}{n-p-1}-1\right)$ goes to zero in probability unconditionally since
${\mathbb{E}}\left(\frac{Z}{n-p-1}\right)=1\;\text{ and }\;\mathrm{Var}\left(\frac{Z}{n-p-1}\right)=\frac{2(n-p-1)}{(n-p-1)^{2}}=\frac{2}{n-p-1},$
so by Chebyshev’s inequality
$\Pr\left(\left|\frac{Z}{n-p-1}-1\right|>\varepsilon\right)\leq\frac{2}{\varepsilon^{2}(n-p-1)}\to 0.$
We previously established that $\frac{n-p-1}{\gamma_{n}}$ converges to 1, and thus the second term in (E.6) satisfies
$\sigma^{2}\left(\frac{Z}{n-p-1}-1\right)\frac{n-p-1}{\gamma_{n}}\overset{P}{\to}0.$
Combining these results, $\sigma^{2}\left(\frac{Z}{n-p-1}-1\right)\frac{n-p-1}{\gamma_{n}}$ and $\sigma^{2}\left(\frac{n-p-1}{\gamma_{n}}-1\right)$ each converge to zero in probability, and hence by the continuous mapping theorem
$\sigma^{2}\left(\frac{Z}{\gamma_{n}}-1\right)=\sigma^{2}\left(\frac{n-p-1}{\gamma_{n}}-1\right)+\sigma^{2}\left(\frac{Z}{n-p-1}-1\right)\frac{n-p-1}{\gamma_{n}}\overset{P}{\to}0.$
By Lemma E.4, unconditional convergence implies conditional convergence, i.e.
$\lim_{n\to\infty}\Pr\left(|{\tilde{\sigma}^{2}}-\sigma^{2}|>\varepsilon\mid Y\in E_{1}\right)=0.$
We can now prove Proposition 2.6.
Proof of Proposition 2.6.
Recall from (10) and (21) that
${p_{H_{0}^{M}\mid E}}:=\Pr\!\Bigl(W/Z\geq r\,\Bigm|\,W-cZ\geq d/\sigma^{2}\Bigr),\quad{p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}:=\Pr\!\Bigl(W/Z\geq r\,\Bigm|\,W-cZ\geq d/{\tilde{\sigma}^{2}}\Bigr),$
where $c$, $d$, and $r$ are defined (as functions of $Y$) in (11) and ${\tilde{\sigma}^{2}}={\tilde{\sigma}^{2}}(0)$ in (20). Note that $c$, $d$, $r$, and $\tilde{\sigma}^{2}$} all depend on $n$. Our goal is to establish that
$\forall\varepsilon>0,\;\;\Pr\bigl(|{p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}-{p_{H_{0}^{M}\mid E}}|>\varepsilon\mid Y\in E_{1}\bigr)\;\to\;0\;\;\text{as}\;\;n\to\infty.$ (E.7)
Define the functions
$N(r^{\prime},c^{\prime},d^{\prime},s^{\prime})=\Pr\Bigl(W/Z\geq r^{\prime},\;W-c^{\prime}Z\geq d^{\prime}/s^{\prime}\Bigr),\quad D(c^{\prime},d^{\prime},s^{\prime})=\Pr\Bigl(W-c^{\prime}Z\geq d^{\prime}/s^{\prime}\Bigr).$
Then
${p_{H_{0}^{M}\mid E}}=\frac{N(r,c,d,\sigma^{2})}{D(c,d,\sigma^{2})},\quad{p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}=\frac{N(r,c,d,{\tilde{\sigma}^{2}})}{D(c,d,{\tilde{\sigma}^{2}})}.$
By Proposition 2.4, $\tilde{\sigma}^{2}$} is a consistent estimator of $\sigma^{2}$. Thus by Slutsky’s theorem, in order to establish that ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}-{p_{H_{0}^{M}\mid E}}$ converges to zero unconditionally, it shall suffice to show that
$N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})\;=\;o_{p}(1),\quad D(c,d,\tilde{\sigma}^{2})-D(c,d,\sigma^{2})\;=\;o_{p}(1).$
Note that a slight variant of Lemma E.3 establishes that $D(c,d,\tilde{\sigma}^{2})$ is bounded away from zero in the limit: since $\tilde{\sigma}^{2}$} is consistent for $\sigma^{2},$ we can redefine $G_{n}(\varepsilon)$ and $G_{n}(\varepsilon,B)$ in the proof of Lemma E.3 to include $|{\tilde{\sigma}^{2}}-\sigma^{2}|<\varepsilon$ so that ${\tilde{\sigma}^{2}}>\sigma^{2}-\varepsilon$, and the rest of the proof follows as before. Therefore $D(c,d,\tilde{\sigma}^{2})$ does not approach zero, so Slutsky’s indeed applies. Once we have unconditional convergence, Lemma E.4 implies convergence of ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}-{p_{H_{0}^{M}\mid E}}$ to zero conditional on $Y\in E_{1}$ (6).
To show that $N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})=o_{p}(1),$ first observe that under $H_{0}^{M}:\beta_{M}=0,$ $r$ in (11) converges to zero as $n\to\infty$. We continue with two cases.
Case 1: Suppose $\beta_{-M}\neq 0$. Similarly to the proof of Lemma E.3, define $\phi:=d/n$ so that $\phi\overset{P}{\to}\phi^{*}$ for a fixed, negative $\phi^{*}<0.$ Define
$E_{n}(\varepsilon)=\{(r,c,\phi,{\tilde{\sigma}^{2}}):r\leq\varepsilon,\;c\leq\varepsilon,\;|\phi-\phi^{*}|\leq\varepsilon,\;|\tilde{\sigma}^{2}-\sigma^{2}|\leq\varepsilon\}.$
For fixed $\eta$ and $\varepsilon$, the following inequalties are satisfied for large enough $n$: $\Pr(r>\varepsilon)<\eta$, $c\leq\varepsilon$, $\Pr(|\phi-\phi^{*}|>\varepsilon)\leq\eta$, and $\Pr(|\tilde{\sigma}^{2}-\sigma^{2}|>\varepsilon)<\eta$ (since $c$ is deterministic and converging to zero, for large enough $n$ we have $c<\varepsilon$ almost surely). Then
$\Pr(E_{n}^{c}(\varepsilon))$ $=\Pr(\{r>\varepsilon\}\;\cup\;\{c>\varepsilon\}\;\cup\;\{|\phi-\phi^{*}|>\varepsilon\}\;\cup\;\{|\tilde{\sigma}^{2}-\sigma^{2}|>\varepsilon\})$
$\leq\Pr(r>\varepsilon)+\Pr(|\phi-\phi^{*}|>\varepsilon)+\Pr(|\tilde{\sigma}^{2}-\sigma^{2}|>\varepsilon)$
$\leq 3\eta.$
Observe that
$\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon\right)$
$=\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon,\;E_{n}(\varepsilon)\right)$
$\hskip 28.45274pt+\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon,\;E_{n}^{c}(\varepsilon)\right)$
$\leq\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon,\;E_{n}(\varepsilon)\right)+\Pr(E_{n}^{c}(\varepsilon)),$
[…]
$(r^{\prime},c^{\prime},\phi^{\prime},s^{\prime})\in[0,\varepsilon]\times[0,\varepsilon]\times[\phi^{*}-\varepsilon,\phi^{*}+\varepsilon]\times[\sigma^{2}-\varepsilon,\sigma^{2}+\varepsilon].$
Thus there exists $\delta>0$ such that $|\tilde{\sigma}^{2}-\sigma^{2}|<\delta$ implies that
$|\tilde{N}(r,c,\phi,\tilde{\sigma}^{2})-\tilde{N}(r,c,\phi,\sigma^{2})|=|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|<\varepsilon$
under the event $E_{n}(\varepsilon)$. This further implies that
$\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon,\;E_{n}(\varepsilon)\right)\leq\Pr\left(|\tilde{\sigma}^{2}-\sigma^{2}|>\delta\right).$
Since $\tilde{\sigma}^{2}$ converges to $\sigma^{2}$ in probability, we have the desired convergence,
$\Pr\left(|N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})|>\varepsilon\right)\to 0,\quad\text{i.e.}\quad N(r,c,d,\tilde{\sigma}^{2})-N(r,c,d,\sigma^{2})=o_{p}(1).$
Case 2: If $\beta=0,$ then $d$ is a scaled, central chi-square random variable with $p$ degrees of freedom. Define the event
$E_{n}(\varepsilon,B)=\{(r,c,d,{\tilde{\sigma}^{2}}):r<\varepsilon,\;c<\varepsilon,\;|d|<B,\;|\tilde{\sigma}^{2}-\sigma^{2}|<\varepsilon\}.$
For fixed $\eta$ and $\varepsilon$, pick $B_{\eta}$ such that for large enough $n$ the following inequalties are satisfied: $\Pr(r>\epsilon)<\eta$, $c<\varepsilon$, $\Pr(|d|>B_{\eta})\leq\eta$, and $\Pr(|\tilde{\sigma}^{2}-\sigma^{2}|>\varepsilon)<\eta$. The rest of the proof of Case 2 follows similarly to the proof of Case 1, where we use pointwise continuity of $N(r^{\prime},c^{\prime},d^{\prime},s^{\prime})$ instead of $\tilde{N}(r^{\prime},c^{\prime},\phi^{\prime},s^{\prime}).$
A similar argument holds to show that $D(c,d,\tilde{\sigma}^{2})-D(c,d,\sigma^{2})=o_{p}(1).$ Therefore $\Pr\bigl(|{p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}-{p_{H_{0}^{M}\mid E}}|>\varepsilon\bigr)\to 0$. Thus, by Lemma E.4, $\Pr\bigl(|{p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}-{p_{H_{0}^{M}\mid E}}|>\varepsilon\mid Y\in E_{1}\bigr)\to 0$. By Theorem 2.1, $p_{H_{0}^{M}\mid E}$} is $\text{Unif}(0,1)$ for all $n$ conditional on $Y\in E_{1}$. It follows that ${p^{\tilde{\sigma}^{2}}_{H_{0}^{M}\mid E}}\mid Y\in E_{1}$ converges in distribution to $\text{Unif}(0,1)$. ∎
E.8 Proof of Proposition 3.1
Proof of Proposition 3.1.
The result follows from algebra, beginning from (10):
${p_{H_{0}^{M}\mid E}}$ $=\Pr\bigg(W/Z\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\Big|W-cZ\geq\frac{1}{\sigma^{2}}\left(-y^{\top}P_{X_{-M}}y\right)\bigg)$
$=\Pr\bigg(W/Z\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}$
$\Big|W-cZ\geq\frac{1}{\sigma^{2}}\left(y^{\top}(P_{X}-P_{X_{-M}})y-y^{\top}P_{X}y\right)\bigg)$
$=\Pr\Bigg(W/Z\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}$
$\Big|W-cZ\geq\frac{1}{\sigma^{2}}\cdot y^{\top}(I_{n-1}-P_{X})y\cdot\left(\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}-\frac{y^{\top}y-y^{\top}(I_{n-1}-P_{X})y}{y^{\top}(I_{n-1}-P_{X})y}\right)\Bigg)$
$=\Pr\Bigg(W/Z\geq\frac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}$
$\Big|W-cZ\geq\tfrac{1}{\sigma^{2}}\tfrac{y^{\top}(I_{n-1}-P_{X})y}{n-p-1}\cdot\bigg(\tfrac{y^{\top}(P_{X}-P_{X_{-M}})y}{y^{\top}(I_{n-1}-P_{X})y}\cdot\tfrac{n-p-1}{m}\cdot m$-
$\hskip 144.54pt\left(1-\tfrac{y^{\top}(I_{n-1}-P_{X})y}{y^{\top}y}\right)(n-p-1)\big(\tfrac{y^{\top}(I_{n-1}-P_{X})y}{y^{\top}y}\big)^{-1}\bigg)\Bigg)$
$=\Pr\Bigg(W/Z\geq F_{H_{0}^{M}}\cdot\frac{m}{n-p-1}$
$\Big|W-cZ\geq\mathrm{RSE}^{2}\cdot\left(F_{H_{0}^{M}}\cdot m-\frac{\mathrm{R}^{2}}{1-\mathrm{R}^{2}}\cdot(n-p-1)\right)/\sigma^{2}\Bigg).$